### v1

# ForAug: Recombining Foregrounds and Backgrounds to Improve Vision Transformer Training with Bias Mitigation

Abstract
Transformers, particularly Vision Transformers (ViTs), have achieved state-of-the-art performance in large-scale image classification. However, they often require large amounts of data and can exhibit biases that limit their robustness and generalizability. This paper introduces ForAug, a novel data augmentation scheme that addresses these challenges and explicitly includes inductive biases, which commonly are part of the neural network architecture, into the training data. ForAug is constructed by using pretrained foundation models to separate and recombine foreground objects with different backgrounds, enabling fine-grained control over image composition during training. It thus increases the data diversity and effective number of training samples. We demonstrate that training on ForNet, the application of ForAug to ImageNet, significantly improves the accuracy of ViTs and other architectures by up to 4.5 percentage points (p.p.) on ImageNet and 7.3 p.p. on downstream tasks. Importantly, ForAug enables novel ways of analyzing model behavior and quantifying biases. Namely, we introduce metrics for background robustness, foreground focus, center bias, and size bias and show that training on ForNet substantially reduces these biases compared to training on ImageNet. In summary, ForAug provides a valuable tool for analyzing and mitigating biases, enabling the development of more robust and reliable computer vision models. Our code and dataset are publicly available at https://github.com/tobna/ForAug.
1 Introduction
Figure 1: Comparison of ForNet and ImageNet. ForNet recombines foreground objects with different backgrounds each epoch, thus creating a more diverse training set. We still apply traditional data augmentation afterwards.
Image classification, a fundamental task in computer vision (CV), involves assigning a label to an image from a predefined set of categories. This seemingly simple task underpins a wide range of applications, including medical diagnosis [39, 50], autonomous driving [52], and object recognition [3, 17, 15]. Furthermore, image classification is used for large-scale pretraining of vision models [10, 31, 47] and to judge the progress of the field of CV [22, 36]. The advent of large-scale datasets, particularly ImageNet [8], containing millions of labeled images across thousands of categories, has been instrumental in driving significant progress in this field. ImageNet served as a catalyst for the rise of large-scale CV models [25, 16] and remains the most important CV benchmark for more than a decade [25, 48, 54, 16].
While traditionally, convolutional neural networks (CNNs) have been the go-to architecture for image classification, Transformers [49], particularly the Vision Transformer (ViT) [10], have emerged as a powerful alternative. These attention-based models have demonstrated superior performance in various vision tasks, including image classification [54, 58, 3, 63, 51].
Data augmentation is a key technique for training image classification models. Traditional data augmentation methods, such as random cropping, flipping, and color jittering, are commonly employed to increase the diversity of the training data and improve the model’s performance [56, 42]. These basic transformations, originally designed for CNNs, change the input images in a way that preserves their semantic meaning [1]. However, the architectural differences of CNNs and Transformers suggest that the latter might benefit from different data augmentation strategies. In particular, the Transformers self-attention mechanism is not translation equivariant [38, 9], meaning that the model does not inherently understand the spatial relationships between pixels.
Inspired by this inductive bias of CNNs, that is not inherent to ViTs, we propose ForAug, a novel data augmentation scheme for image classification which makes the translation equivariance of CNNs explicit in the training data by recombining foreground objects at varying positions with different backgrounds. Applying ForAug to ImageNet gives rise to ForNet, a novel dataset that enables this data augmentation with with fine-grained control over the image composition. Recognizing that Transformers need to learn the spatial relationships from data, since they are not inherently translation invariant, and in general are usually trained on larger datasets [24], we separate the foreground objects in ImageNet from their backgrounds, using an open-world object detector [37], and fill in the background in a plausible way using an object removal model [43, 45]. This allows us to recombine any foreground object with any background on the fly, creating a highly diverse training set. During recombination, we can control important parameters, like the size and position of the foreground object, to help the model learn the spatial invariances necessary for image classification. We show that training on ForNet instead of ImageNet increases the model accuracy of Transformers by up to 4.5 p.p. on ImageNet and an up to $39.3\%$ reduction in error rate on downstream tasks.
Additionally, ForAug is a useful tool for analyzing model behavior and biases, when used during the evaluation phase. We utilize our control over the image distribution to quantify a model’s background robustness (by varying the choice of background), foreground focus (by leveraging our knowledge about the placement of the foreground object), center bias (by controlling the object’s position), and size bias (by controlling object size). These analyses provide insights into model behavior and biases, which is crucial for model deployment and future robustness optimizations. We show that training on ForNet, instead of ImageNet, significantly reduces all of these biases, completely removing the models’ dependence on the background distribution. We make our code for ForAug and the ForNet-dataset publicly available [^1] https://github.com/tobna/ForAug to facilitate further research.
Contributions
  • We propose ForAug, a novel data augmentation scheme, that recombines objects and backgrounds to train Transformers for image classification.
  • We show that training on ForNet, the ImageNet instantiation of ForAug, leads to 4.5 p.p. improved accuracy on ImageNet and 7.3 p.p. on downstream tasks.
  • We propose novel ForAug-based metrics to analyze and quantify fine-grained biases trained models: Background Robustness, Foreground Focus, Center Bias, and Size Bias. Training on ForNet, instead of ImageNet, significantly reduces these biases.
2 Related Work
Data Augmentation for Image Classification
Data augmentation is a crucial technique for improving the performance and generalization of image classification models. Traditional augmentation strategies rely on simple geometric or color-space transformations like cropping, flipping, roatation, blurring, color jittering, or random erasing [62] to increase the diversity of the training data without changing their semantic meaning. With the advent of Transformers, new data augmentation operations like PatchDropout [30] have been proposed. Other transformations like Mixup [61], CutMix [59], or random cropping and patching [46] combine multiple input images. These simple transformations are usually bundled to form more complex augmentation policies like AutoAugment [5] and RandAugment [6], which automatically search for optimal augmentation policies or 3-augment [48] which is optimized to train a ViT. For a general overview of data augmentation techniques for image classification, we refer to [42, 56].
We build upon these general augmentation techniques by introducing a novel approach to explicitly separate and recombine foregrounds and backgrounds for image classification. Our approach is used in tandem with traditional data augmentation techniques to improve model performance and reduce biases.
Copy-Paste Augmentation
The copy-paste augmentation [14], which is used for object detection [41, 14] and instance segmentation [53, 28], involves copying segmented objects from one image and pasting them onto another. While typically human-annotated segmentation masks are used to extract the foreground objects, other foregound sources have been explored, like 3D models [19] and pretrained object-detection models for use on objects on white background [11] or synthetic images [12]. DeePaste [53] focuses on using inpainting for a more seamless integration of the pasted object.
Unlike these methods, ForNet focuses on image classification. While for detection and segmentation, objects are pasted onto another image (with a different foreground) or on available or rendered background images of the target scene, we extract foreground objects and fill in the resulting holes in the background in a semantically neutral way. This way, we can recombine any foreground object with a large variety of neutral backgrounds from natural images, enabling a controlled and diverse manipulation of image composition.
Model robustness evaluation
Evaluating model robustness to various image variations is critical for understanding and improving model generalization. Datasets like ImageNet-C [18] and ImageNet-P [18] introduce common corruptions and perturbations. ImageNet-E [27] evaluates model robustness against a collection of distribution shifts. Other datasets, such as ImageNet-D [60], focus on varying background, texture, and material, but rely on synthetic data. Stylized ImageNet [13] investigates the impact of texture changes. ImageNet-9 [55] explores background variations using segmented images, but the backgrounds are often artificial.
In contrast to these existing datasets, which are used only for evaluation, ForNet provides fine-grained control over foreground object placement, size, and background selection, enabling a precise and comprehensive analysis of specific model biases within the context of a large-scale, real-world image distribution. As ForNet also provides controllable training set generation, it goes beyond simply measuring robustness to actively improving it through training.
3 RecombiNet (Method)
Figure 2: Overview of ForNet. The data creation consists of two stages: (1, offline) Segmentation, where we segment the foreground objects from the background and fill in the background. (2, online) Recombination, where we combine the foreground objects with different backgrounds to create new samples.
We introduce ForAug, a data augmentation scheme designed to enhance Transformer training by explicitly separating and recombining foreground objects and backgrounds. ForAug involves two stages: Segmentation and Recombination, both visualized in Figure 2.
Segmentation
The segmentation stage isolates the foreground objects and their corresponding backgrounds. We then fill in the background in a visually plausible way [43] using a pretrained object-removal model. This stage is computed once offline and the results are stored for the recombination stage.
First, foreground objects are detected and segmented from their backgrounds using a prompt-based segmentation model to exploit the classification datasets labels. We use the state-of-the-art Grounded SAM [37], which is based on Grounding DINO [29] and SAM [23]. The prompt we use is “a <class name>, a type of <object category>”, where <class name> is the specific name of the objects class as defined by the dataset and <object category> is a the broader category of the object. The <object category> guides the segmentation model towards the correct object in case the <class name> alone is too specific. This can be the case with prompts like “sorrel” or “guenon”, where the more general name “horse” or “monkey” is more helpful. We derive the <object category> from the WordNet hierarchy, using the immediate hypernym.
We iteratively extract up to $n$ foreground masks for each dataset-image, using different more and more general prompts based on the more general synsets of WordNet (e.g. “a sorrel, a type of horse”, “a horse, a type of equine”, …). Masks that are very similar, with a pairwise IoU of at least $0.9$, are merged. The output is a set of masks delineating the foreground objects and the backgrounds. We select the best mask per image (according to Equation 1) in a later filtering step, described below.
An inpainting model that is specifically optimized to remove objects from images, such as LaMa [45] or Attentive Eraser [43], is used to inpaint the foreground regions in the backgrounds. To ensure the quality of the foreground and background images (for each dataset-image), we select a foreground/background pair from the $\leq n$ variants we have extracted and infilled in the previous steps. Using an ensemble of six ViT, ResNet, and Swin Transformer models pretrained on the original dataset, we select the foreground/background pair that maximizes foreground performance while minimizing the performance on the background and size of the foreground according to:
$\begin{split}\text{score}(\mathrm{fg},\mathrm{bg},c)&=\log\left(\frac{1}{\left|E\right|}\sum_{m\in E}\mathbb{P}[m(\mathrm{fg})=c]\right)\\ &+\log\left(1-\frac{1}{\left|E\right|}\sum_{m\in E}\mathbb{P}[m(\mathrm{bg})=c]\right)\\ &+\lambda\log\left(1-\left|\frac{\operatorname{size}(\mathrm{fg})}{\operatorname{size}(\mathrm{bg})}-\varepsilon\right|\right).\end{split}$ (1)
Here, $E$ is the ensemble of models and $m$ is a pretrained model, $c$ is the correct foreground class, $\mathrm{fg}$, and $\mathrm{bg}$ are the foreground and background and $\operatorname{size}(\cdot)$ is the size in number of pixels. We ran a hyperparameter search using a manually annotated subset of foreground/background variants to find the factors in Equation 1: $\lambda=2$ and $\varepsilon=0.1$. The optimal foreground size of $10\%$ of the full image balances the smallest possible foreground size that encompasses all the respective class information in the image with still conveying the foreground information after pasting it onto another background. This filtering step ensures we segment all the relevant foreground objects.
Finally, we filter out backgrounds that are more than $80\%$ infilled, as these tend to be overly synthetic, plain and don’t carry much information (see Appendix C). We ablate this choice in Section 4.1. In summary, we factorize the dataset into a set of foreground objects with a transparent background and a set of diverse backgrounds per class. The next step is to recombine them as data augmentation before applying common data augmentation operations during training.
Recombination
The recombination stage, which is performed online, combines the foreground objects with different backgrounds to create new training samples. For each object, we follow the pipeline of: Pick an appropriate background, resize it to a fitting size, place it in the background image, smooth the transition edge, and apply other data augmentations.
For each foreground object, we sample a background using one of the following strategies: (1) the original image background, (2) the set of backgrounds from the same class, or (3) the set of all possible backgrounds. These sets are trading off the amount of information the model can learn from the background against the diversity of new images created. In each epoch, each foreground object is seen exactly once, but a background may appear multiple times.
The selected foreground is resized based on its relative size within its original image and the relative size of the original foreground in the selected background image. The final size is randomly selected from a 30% range around upper and lower limits ($s_{u}$ and $s_{l}$), based on the original sizes:
$s\sim\mathcal{U}\left[(1-0.3)s_{l},(1+0.3)s_{u}\right].$ (2)
To balance the size of the foreground and that of the backgrounds original foreground, the upper and lower limit $s_{u}$ and $s_{l}$ are set to the mean or range of both sizes, depending on the foreground size strategy: mean or range.
The resized foreground is then placed at a random position within the background image. This position is sampled from a generalization of the Bates distribution [2] with parameter $\eta\in\mathbb{N}$, visualized in Figure 3. We choose the bates distribution, as it presents an easy way to sample from a bounded domain with just one hyperparameter that controls the concentration of the distribution. $\eta=1$ corresponds to the uniform distribution; $\eta>1$ concentrates the distribution around the center; and for $\eta<-1$, the distribution is concentrated at the borders. To more seamlessly integrate the foreground, we apply a Gaussian blur with ${\sigma\in[\frac{\sigma_{\text{max}}}{10},\sigma_{\text{max}}]}$, inspired by the standard range for the Gaussian blur operation in [48], to the foreground’s alpha-mask.
We can apply standard data augmentation techniques in two modes: Either we apply all augmentations to the recombined image, or we apply the cropping and resizing to the background only and then apply the other augmentations after recombination. The second mode ensures the foreground object remains fully visible, while the first mode mirrors standard data augmentation practices.
We experiment with a constant mixing ratio, or a linear or cosine annealing schedule that increases the amount of images from the original dataset over time. The mixing ratio acts as a probability of selecting an image from the original dataset; otherwise, an image with the same foreground is recombined using ForAug. Thus, we still ensure each foreground is seen once per epoch.
4 Experiments
We conduct a comprehensive suit of experiments to validate the effectiveness of our approach. We compare training on ForNet, the ImageNet instantiation of ForAug, to training on ImageNet for 7 different models. Furthermore, we assess the impact of using ForNet for pretraining on multiple fine-grained downstream datasets. Additionally, we use ForAug’s control over the image distribution to quantify some model behaviors and biases.
4.1 Design Choices of ForAug
We start by ablating the design choices of ForAug. For this, we revert to TinyImageNet [26], a subset of ImageNet containing 200 categories with 500 images each, and TinyForNet, a version of ForAug derived from TinyImageNet. Table 1 presents the results of these ablations.
Dataset Detect. Infill FG. Augmentation BG. BG. edge original TinyImageNet Accuracy
prompt Model size Order strategy pruning smoothing image mixing ViT-Ti [%] ViT-S [%]
TinyImageNet $66.1\pm 0.5$ $68.3\pm 0.7$
[…]
Table 1: Ablation of design decisions of TinyForNet on TinyImageNet and ForNet on ImageNet.
Prompt. First, we evaluate the type of prompt used to detect the foreground object. Here, the general prompt, which contains the class and the more general object category, outperforms only having the class name (specific).
Inpainting. Attentive Eraser [43] produces superior results compared to LaMa [45] (see Appendix B for examples).
Foreground size significantly impacts performance. Employing a range of sizes during recombination, rather than a fixed mean size, boosts accuracy by approximately 1 p.p. This suggests that the added variability is beneficial.
[…]
Background pruning. When it comes to the choice of backgrounds to use, we test two pruning thresholds ($t_{\text{prune}}$) to exclude backgrounds with excessive inpainting. A threshold of $t_{\text{prune}}=1.0$ means that we use all backgrounds that are not fully infilled. Varying $t_{\text{prune}}$ has minimal impact. Therefore, we choose $t_{\text{prune}}=0.8$ to exclude predominantly artificial backgrounds. Similarly, applying edge smoothing to foreground masks with Gaussian blurring actually hurts performance on TinyForNet, but slightly improves it on ForNet.
Mixing ForNet with the original ImageNet data proves crucial. While constant and linear mixing schedules improve performance over no mixing by $2-3$ p.p. compared to only using TinyForNet, the cosine annealing schedule yields the best results, boosting accuracy by another $0.5-1$ p.p.
Background strategy. Another point is the allowed choice of background image for each foreground object. We compare using the original background, a background from the same class, and any background. These strategies go from low diversity and high shared information content between the foreground and background to high diversity and low shared information content. For ViT-Ti, the latter two strategies perform comparably, while ViT-S benefits from the added diversity of using any background. The same is true when training on the full (ImageNet) version of ForNet.
Figure 3: Plot of the probability distribution function (PDF) of the extended Bates distribution for different parameters $\eta$. Higher values of $\eta$ concentrate the distribution around the center.
Training Set/ Bates Parameter TIN TinyForNet
$\eta=-3$ $-2$ $/-1$ 1 $2$ $3$
TinyImageNet 68.9 60.5 60.2 60.8 62.6 63.1
$\eta=-3$ 71.3 79.3 79.5 79.1 79.3 79.1
$\eta=-2$ 71.5 80.0 78.7 79.3 79.1 78.8
$\eta=1/-1$ 72.3 79.5 78.9 80.2 79.7 80.4
$\eta=2$ 71.3 78.2 77.8 79.1 79.6 79.9
$\eta=3$ 71.4 77.2 76.9 78.6 79.6 79.7
Table 2: Accuracy of ViT-S trained on TinyImageNet (TIN) and TinyForNet with different foreground position distributions by varying the parameter of a Bates distribution $\eta$. The best performance is achieved using the uniform distribution ($\eta=1$).
Foreground position. Finally, we analyze the foreground object’s positioning in the image. We utilize an extended Bates distribution to sample the position of the foreground object. The Bates distribution [2] with parameter $\eta\geq 1$ is the mean of $\eta$ independent uniformly distributed random variables [20]. Therefore, the larger $\eta$, the more concentrated the distribution is around the center. We extend this concept to $\eta\leq-1$ by defining ${X\sim\text{Bates}(\eta):\Leftrightarrow s(X)\sim\text{Bates}(-\eta)}$ for $\eta\leq 1$ with $s$ being the sawtooth function on $[0,1]$:
$s(x)=\begin{cases}x+0.5&\text{if }0<x<0.5\\ x-0.5&\text{if }0.5\leq x\leq 1\end{cases}$ (3)
Note that $s\circ s=\mathrm{id}$ on $[0,1]$. This way, distributions with $\eta\leq-1$ are more concentrated around the borders. $\eta=1$ and $\eta=-1$ both correspond to the uniform distribution. The PDF of this extended Bates distribution is visualized in Figure 3.
When sampling more towards the center of the image, the difficulty of the task is reduced, which then reduces the performance on TinyImageNet. This is reflected in the performance when evaluating on TinyForNet with $\eta=2$ and $\eta=3$ compared to $\eta=-1/1$. We observe a similar reduction for $\eta<-1$. This experiment is conducted using the LaMa infill model.
Dataset Classes Training Images Validation Images
TinyImageNet 200 100,000 10,000
TinyForNet 200 99,404 9,915
ImageNet 1,000 1,281,167 50,000
ForNet 1,000 1,274,557 49,751
Table 3: Dataset statistics for TinyImageNet, TinyForNet, ImageNet, and ForNet. For ForNet and TinyForNet we report the number of foreground/background pairs.
After fixing the optimal design parameters in Table 1 (last row), we construct the full ForNet dataset using the entire ImageNet dataset. Table 3 compares the dataset statistics of ImageNet and ForNet. The slightly reduced image count in ForNet is due to instances where Grounded SAM failed to produce valid object detections.
4.2 Image Classification Results
Model ImageNet Accuracy when trained on Delta
ImageNet ForNet
ViT-S $79.1\pm 0.1$ $81.4\pm 0.1$ +2.3
ViT-B $77.6\pm 0.2$ $81.1\pm 0.4$ +3.5
ViT-L $75.3\pm 0.4$ $79.8\pm 0.1$ +4.5
Swin-Ti $77.9\pm 0.2$ $79.7\pm 0.1$ +1.8
Swin-S $79.4\pm 0.1$ $80.6\pm 0.1$ +1.2
ResNet-50 $78.3\pm 0.1$ $78.8\pm 0.1$ +0.5
ResNet-101 $79.4\pm 0.1$ $80.4\pm 0.1$ +1.0
Table 4: ImageNet results of models trained on ForNet and on ImageNet directly. ForNet improves the performance of all models in our test.
Table 4 compares the ImageNet performance of models trained on ForNet and ones trained directly on ImageNet. We adopt the training setup of [33] and [48] (details in Appendix A) for training ViT [10], Swin [31] and ResNet [16] models. Notably, ForNet improves performance across all tested architectures, including the ResNet models (up to $1$ p.p.), demonstrating benefits beyond Transformers. For Transformer models, we observe improvements from $1.2$ p.p. to $4.5$ p.p. This improvement is more substantial for the larger models, with ViT-L gaining $4.5$ p.p. in accuracy. ForNet’s improvements mostly counteract the drop in performance due to overfitting for large models. When training on ImageNet, this drop is $3.8$ p.p. from ViT-S to ViT-L, while for ForNet it is reduced to $1.6$ p.p.
Model Aircraft Cars Flowers Food Pets
ViT-S @ ImageNet $72.4\pm 1.0$ $89.8\pm 0.3$ $94.5\pm 0.2$ $89.1\pm 0.1$ $93.8\pm 0.2$
ViT-S @ ForNet $78.6\pm 0.5$ $92.2\pm 0.2$ $95.5\pm 0.2$ $89.6\pm 0.1$ $94.5\pm 0.2$
+6.2 +2.4 +1.0 +0.5 +0.7
ViT-B @ ImageNet $71.7\pm 0.5$ $90.0\pm 0.2$ $94.8\pm 0.4$ $89.8\pm 0.2$ $94.1\pm 0.4$
ViT-B @ ForNet $79.0\pm 2.2$ $93.3\pm 0.1$ $96.5\pm 0.1$ $90.9\pm 0.1$ $95.1\pm 0.4$
+7.3 +3.3 +1.7 +1.1 +1.0
ViT-L @ ImageNet $72.1\pm 1.0$ $88.8\pm 0.3$ $94.4\pm 0.3$ $90.1\pm 0.2$ $94.2\pm 0.4$
ViT-L @ ForNet $77.6\pm 1.2$ $89.1\pm 0.2$ $96.6\pm 0.1$ $91.3\pm 0.1$ $95.1\pm 0.1$
+5.5 +0.3 +2.2 +1.2 +0.9
Swin-Ti @ ImageNet $77.0\pm 0.1$ $91.3\pm 0.6$ $95.9\pm 0.1$ $90.0\pm 0.2$ $94.2\pm 0.1$
Swin-Ti @ ForNet $81.1\pm 0.8$ $92.8\pm 0.4$ $96.2\pm 0.1$ $90.4\pm 0.3$ $94.8\pm 0.5$
+4.1 +2.5 +0.3 +0.4 +0.6
Swin-S @ ImageNet $75.7\pm 1.4$ $91.0\pm 0.3$ $95.9\pm 0.5$ $91.1\pm 0.2$ $94.4\pm 0.1$
Swin-S @ ForNet $81.4\pm 0.2$ $93.1\pm 0.2$ $96.3\pm 0.3$ $91.2\pm 0.2$ $94.9\pm 0.3$
+5.7 +2.1 +1.4 +0.1 +0.5
ResNet-50 @ ImageNet $78.2\pm 0.5$ $89.8\pm 0.2$ $91.7\pm 0.4$ $84.4\pm 0.2$ $93.7\pm 0.3$
ResNet-50 @ ForNet $80.3\pm 0.4$ $90.4\pm 0.2$ $91.7\pm 0.2$ $84.5\pm 0.2$ $93.7\pm 0.3$
+2.1 +0.6 $\pm$0 +0.1 $\pm$0
ResNet-101 @ ImageNet $78.4\pm 0.6$ $90.3\pm 0.1$ $91.2\pm 0.5$ $86.0\pm 0.2$ $94.3\pm 0.2$
ResNet-101 @ ForNet $81.4\pm 0.5$ $91.3\pm 0.1$ $92.9\pm 0.2$ $86.3\pm 0.1$ $94.0\pm 0.3$
+3.0 +1.3 +1.7 +0.3 -0.3
Table 5: Downstream accuracy in percent when finetuning on other datasets. Models were pretrained on ForNet and ImageNet. Pretraining on ForNet increases Transformer downstream accuracy on all datasets.
To assess the transferability of ForNet-trained models, we finetune models pretrained on ImageNet and ForNet on five fine-grained datasets: FGVC-Aircraft [32], Stanford Cars [7], Oxford Flowers [34], Food-101 [21], and Oxford-IIIT Pets [35]. While for ResNets, the performance of both training datasets is about the same, for every Transformer, we see the accuracy improve on all downstream dataset by up to 7.3 p.p. and a reduction of error rate of up to $39.3\%$. In summary, these results demonstrate that the improved representation learning achieved by training on ForNet translates to superior performance not only on ImageNet, but also on a variety of fine-grained image classification tasks.
4.3 Further Model Evaluation
Beyond its use for training, ForNet’s unique properties and controlled data generation capabilities make it a powerful tool for analyzing model behavior and biases.
Background Robustness
Model Background Robustness when trained on Delta
ImageNet ForNet
ViT-S $0.73\pm 0.01$ $0.99\pm 0.01$ +0.26
ViT-B $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
ViT-L $0.70\pm 0.01$ $1.00\pm 0.01$ +0.30
Swin-Ti $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
Swin-S $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
ResNet-50 $0.79\pm 0.01$ $0.99\pm 0.01$ +0.20
ResNet-101 $0.79\pm 0.01$ $1.00\pm 0.01$ +0.21
Table 6: Evaluation of the background robustness of models trained on ForNet and on ImageNet directly. Training on ForNet improves the background robustness of all model to $\approx 1.00$, meaning the model is indifferent to the choice of background.
We assess the robustness of models to shifts in the background distribution from a class-related background to any background. Background robustness is defined to be the ratio of accuracy on ForNet with same-class backgrounds to accuracy with any background:
Background $\text{Background Robustness}=\frac{\text{Acc}(\textit{ForNet}_{\text{all}})}{\text{Acc}(\textit{ForNet}_{\text{same}})}$ (4)
It represents the relative drop in performance under a background distribution shift. Table 6 presents the background robustness of various models. When trained on ImageNet, smaller models generally exhibit greater robustness to changes in the background distribution than larger models and ResNet is more robust than the tested Transformer models. Crucially, training on ForNet instead of ImageNet improves the background robustness of all models to $\approx 1.00$, meaning that these models are agnostic to the choice of background and only classify based on the foreground. These findings highlight the generalization benefits of ForNet.
Foreground Focus
Model Foreground Focus when trained on
IN FN IN FN IN FN
GradCam GradCam++ IG
ViT-S $1.2\pm 0.1$ $2.3\pm 0.3$ $1.2\pm 0.1$ $2.1\pm 0.4$ $1.9\pm 0.1$ $2.7\pm 0.1$
ViT-B $1.2\pm 0.1$ $2.4\pm 0.7$ $1.1\pm 0.1$ $2.1\pm 0.1$ $1.7\pm 0.1$ $2.7\pm 0.1$
ViT-L $1.3\pm 0.1$ $1.6\pm 0.1$ $1.1\pm 0.1$ $1.3\pm 0.1$ $1.3\pm 0.1$ $2.6\pm 0.1$
Swin-Ti $0.9\pm 0.1$ $0.7\pm 0.1$ $1.0\pm 0.3$ $0.7\pm 0.3$ $2.5\pm 01$ $4.8\pm 0.3$
Swin-S $0.8\pm 0.1$ $0.7\pm 0.1$ $0.7\pm 0.1$ $0.7\pm 0.4$ $2.4\pm 0.1$ $4.6\pm 0.3$
ResNet-50 $2.2\pm 0.1$ $2.7\pm 0.1$ $2.0\pm 0.1$ $2.9\pm 0.1$ $3.2\pm 0.1$ $4.9\pm 0.2$
ResNet-101 $2.3\pm 0.1$ $2.8\pm 0.1$ $2.2\pm 0.1$ $3.0\pm 0.1$ $3.2\pm 0.1$ $4.8\pm 0.1$
Table 7: Evaluation of the foreground focus using GradCam, GradCam++ and IntegratedGradients of models trained on ForNet (FN) and on ImageNet (IN) directly. Training on ForNet improves the foreground focus of almost all models.
Leveraging our inherent knowledge of the foreground masks when using ForNet, as well as common XAI techniques [40, 4, 44], we can evaluate a model’s focus on the foreground object. We can directly evaluate ImageNet trained models, but this technique can also be extended to other datasets without relying on manually annotated foreground-masks. To evaluate the foreground focus, we employ Grad-CAM [40], Grad-CAM++ [4] or IntegratedGradients (IG) [44] to compute the per-pixel importance of an image for the model’s prediction. The foreground focus is defined to be the ratio of the foreground’s relative importance to its relative size in the image:
[…]
The foreground focus of a model is its average foreground focus over all test images. Table 7 presents our findings. Training on ForNet significantly increasees the foreground focus of ViT and ResNet across all metrics used. For Swin, the foreground focus stagnates when measured using GradCam and GradCam++, but almost doubles when using IG.
Center Bias
Model Center Bias when trained on Delta
ImageNet ForNet
ViT-S [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.255\pm 0.008$ $0.220\pm 003$0.035
ViT-B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.254\pm 0.004$ $0.190\pm 0.002$0.064
ViT-L [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.243\pm 0.011$ $0.117\pm 0.007$0.126
Swin-Ti [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.250\pm 0.007$ $0.165\pm 0.002$0.085
Swin-S [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.232\pm 0.001$ $0.156\pm 002$0.076
ResNet50 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.263\pm 0.003$ $0.197\pm 0.003$0.066
ResNet101 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.230\pm 0.003$ $0.199\pm 002$0.031
Table 8: Evaluation of the position bias. We plot the accuracy relative to the center accuracy of multiple instantiations of the models when the foreground objects is in different cells a $3\times 3$ grid. Training on ForNet significantly reduces a models center bias.
With ForNet we have unique control over the position of the foreground object in the image. This lets us quantify the center bias of ImageNet- and ForNet-trained models. We divide the image into a $3\times 3$ grid and evaluate model accuracy when the foreground object is in each of the $9$ grid cells. Each cell’s accuracy is divided by the accuracy in the center cell for normalization, which gives us the relative performance drop when the foreground is in each part of the image. The center bias is calculated as one minus the average of the minimum performance of a corner cell and the minimum performance of a side cell:
Center $\begin{split}&\text{Center Bias}=\\ &\hskip 7.0pt1-\frac{\min\limits_{a,b\in\{0,2\}}\text{Acc}(\text{cell}_{(a,b)})+\min\limits_{\begin{subarray}{c}a=1\text{ or }b=1\\ a\neq b\end{subarray}}\text{Acc}(\text{cell}_{(a,b)})}{2\text{Acc}(\text{cell}_{(1,1)})}\end{split}$ (6)
Table 8 visualizes the center bias of three instantiations of each model. Performance is generally highest in the center and the center top and bottom and center left and right cells, and lowest in the four corners. Interestingly, ImageNet-trained models perform slightly better when the foreground object is on the right side of the image, compared to the left side, despite our use of random flipping with a probability of $0.5$ during training. Training on ForNet significantly reduces center bias across all models. This demonstrates that ForNet promotes a more uniform spatial attention distribution. Their accuracy is higher in the center left and right cells than in the center top and bottom ones, which is not the case for ImageNet-trained models.
Size Bias
Figure 4: Evaluation of the size bias of models trained on ForNet. We plot the accuracy relative to the accuracy when using the mean foreground size.
Finally, we evaluate the impact of different-sized foreground objects on the accuracy. For this evaluation, we use the mean foreground size strategy. We introduce a size factor $f_{\text{size}}$ by which we additionally scale the foreground object before pasting it onto the background. Results are again normalized by the accuracy when using the mean foreground size ($f_{\text{size}}=1.0$). Figure 4 shows the size bias curves of ViT-S and ViT-B when trained on ImageNet and ForNet. Models trained on ForNet maintain better performance even with smaller foreground objects, when ImageNet-trained models exhibit a more rapid performance decline. Therefore, ForNet-training improves robustness to variations in object scale.
5 Discussion & Conclusion
We introduce ForAug, a novel data augmentation scheme that facilitates improved Transformer training for image classification. By explicitly separating and recombining foreground objects and backgrounds, ForAug enables controlled data augmentation, leading to significant performance gains on ImageNet and downstream fine-grained classification tasks. Furthermore, ForAug provides a powerful framework for analyzing model behavior and quantifying biases, including background robustness, foreground focus, center bias, and size bias. Our experiments demonstrate that training on ForNet, the instantiation of ForAug on ImageNet, not only boosts accuracy but also significantly reduces these biases, resulting in more robust and generalizable models. In the future, we see ForAug be also applied to other datasets and tasks, like video recognition or segmentation. ForAug’s ability to both improve performance and provide insights into model behavior makes it a valuable tool for advancing CV research and developing more reliable AI systems.
Acknowledgements
This work was funded by the Carl-Zeiss Foundation under the Sustainable Embedded AI project (P2021-02-009) and by the EU project SustainML (Horizon Europe grant agreement No 101070408). All compute was done thanks to the Pegasus cluster at DFKI.
Appendix A Training Setup
Parameter Value
Image Resolution $224\times 224$
Epochs 300
Learning Rate 3e-3
Learning Rate Schedule cosine decay
Batch Size 2048
Warmup Schedule linear
Warmup Epochs 3
Weight Decay 0.02
Label Smoothing 0.1
Optimizer Lamb [57]
Data Augmentation Policy 3-Augment [48]
Table 9: Training setup for our ImageNet and ForNet training.
Dataset Batch Size Epochs Learning Rate
Aircraft 512 500 3e-4
Cars 1024 500 3e-4
Flowers 256 500 3e-4
Food 2048 100 3e-4
Pets 512 500 3e-4
Table 10: Training setup for finetuning on different downstream datasets. Other settings are the same as in Table 9.
On ImageNet we use the same training setup as [33] and [48] without pretraining. As our focus is on evaluating the changes in accuracy due to ForAug/ForNet, like [33], we stick to one set of hyperparameters for all models. We list the settings used for training on ImageNet and ForNet in Table 9 and the ones used for finetuning those weights on the downstream datasets in Table 10.
Appendix B Infill Model Comparison
LaMa Att. Eraser LaMa Att. Eraser
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 11: Example infills of LaMa and Attentive Eraser.
Appendix C Images with High Infill Ratio
Infill Ratio LaMa Att. Eraser
93.7 [Uncaptioned image] [Uncaptioned image]
95.7 [Uncaptioned image] [Uncaptioned image]
83.7 [Uncaptioned image] [Uncaptioned image]
88.2 [Uncaptioned image] [Uncaptioned image]
Table 12: Example infills with a large relative foreground area size that is infilled (infill ratio).

### v2

@@
-ForAug: Recombining Foregrounds and Backgrounds to Improve Vision Transformer Training with Bias Mitigation
+ForAug: Mitigating Biases and Improving Vision Transformer Training by Recombining Foregrounds and Backgrounds
 Abstract
 Transformers, particularly Vision Transformers (ViTs), have achieved state-of-the-art performance in large-scale image classification.
-However, they often require large amounts of data and can exhibit biases that limit their robustness and generalizability.
-This paper introduces ForAug, a novel data augmentation scheme that addresses these challenges and explicitly includes inductive biases, which commonly are part of the neural network architecture, into the training data.
-ForAug is constructed by using pretrained foundation models to separate and recombine foreground objects with different backgrounds, enabling fine-grained control over image composition during training.
-It thus increases the data diversity and effective number of training samples.
-We demonstrate that training on ForNet, the application of ForAug to ImageNet, significantly improves the accuracy of ViTs and other architectures by up to 4.5 percentage points (p.p.) on ImageNet and 7.3 p.p.
+However, they often require large amounts of data and can exhibit biases, such as center or size bias, that limit their robustness and generalizability.
+This paper introduces ForAug, a novel data augmentation operation that addresses these challenges by explicitly imposing invariances into the training data, which are otherwise part of the neural network architecture.
+ForAug is constructed by using pretrained foundation models to separate and recombine foreground objects with different backgrounds.
+This recombination step enables us to take fine-grained control over object position and size, as well as background selection.
+We demonstrate that using ForAug significantly improves the accuracy of ViTs and other architectures by up to 4.5 percentage points (p.p.) on ImageNet, which translates to 7.3 p.p.
 on downstream tasks.
-Importantly, ForAug enables novel ways of analyzing model behavior and quantifying biases.
-Namely, we introduce metrics for background robustness, foreground focus, center bias, and size bias and show that training on ForNet substantially reduces these biases compared to training on ImageNet.
+Importantly, ForAug not only improves accuracy but also opens new ways to analyze model behavior and quantify biases.
+Namely, we introduce metrics for background robustness, foreground focus, center bias, and size bias and show that using ForAug during training substantially reduces these biases.
 In summary, ForAug provides a valuable tool for analyzing and mitigating biases, enabling the development of more robust and reliable computer vision models.
-Our code and dataset are publicly available at https://github.com/tobna/ForAug.
+Our code and dataset are publicly available at <url>.
 1 Introduction
-Figure 1: Comparison of ForNet and ImageNet.
-ForNet recombines foreground objects with different backgrounds each epoch, thus creating a more diverse training set.
-We still apply traditional data augmentation afterwards.
-Image classification, a fundamental task in computer vision (CV), involves assigning a label to an image from a predefined set of categories.
-This seemingly simple task underpins a wide range of applications, including medical diagnosis [39, 50], autonomous driving [52], and object recognition [3, 17, 15].
-Furthermore, image classification is used for large-scale pretraining of vision models [10, 31, 47] and to judge the progress of the field of CV [22, 36].
-The advent of large-scale datasets, particularly ImageNet [8], containing millions of labeled images across thousands of categories, has been instrumental in driving significant progress in this field.
-ImageNet served as a catalyst for the rise of large-scale CV models [25, 16] and remains the most important CV benchmark for more than a decade [25, 48, 54, 16].
-While traditionally, convolutional neural networks (CNNs) have been the go-to architecture for image classification, Transformers [49], particularly the Vision Transformer (ViT) [10], have emerged as a powerful alternative.
-These attention-based models have demonstrated superior performance in various vision tasks, including image classification [54, 58, 3, 63, 51].
+Figure 1: Comparison of traditional image classification training and training when using ForAug.
+ForAug recombines foreground objects with different backgrounds each epoch, thus creating a more diverse training set.
+We still apply strong traditional data augmentation afterwards.
+Image classification, a fundamental task in computer vision (CV), involves assigning labels to images from a set of categories.
+It underpins a wide range of applications, like medical diagnosis [39, 50], autonomous driving [53], and object recognition [2, 16, 14] and facilitates large-scale pretraining [9, 30, 47], and progress evaluation in CV [22, 36].
+The advent of large-scale datasets, particularly ImageNet [7], served as a catalyst for the rise of large-scale CV models [24, 15] and remains the most important CV benchmark for more than a decade [24, 48, 56, 15].
+While traditionally, convolutional neural networks (CNNs) have been the go-to architecture in CV, Transformers [49], particularly the Vision Transformer (ViT) [9], have emerged as a powerful alternative and go-to architecture, demonstrating superior performance in various vision tasks, including image classification [56, 60, 2, 65, 52].
 Data augmentation is a key technique for training image classification models.
-Traditional data augmentation methods, such as random cropping, flipping, and color jittering, are commonly employed to increase the diversity of the training data and improve the model’s performance [56, 42].
-These basic transformations, originally designed for CNNs, change the input images in a way that preserves their semantic meaning [1].
+Traditional augmentation methods, such as cropping, flipping, or color shifts, are commonly employed to increase data diversity [58, 42], but remain bound to existing image compositions.
+While these preserve the images’ semantic meaning, their ability to teach spatial invariances is limited.
+While combinations of these data augmentations are still used today, they originally were proposed to benefit CNNs.
 However, the architectural differences of CNNs and Transformers suggest that the latter might benefit from different data augmentation strategies.
-In particular, the Transformers self-attention mechanism is not translation equivariant [38, 9], meaning that the model does not inherently understand the spatial relationships between pixels.
-Inspired by this inductive bias of CNNs, that is not inherent to ViTs, we propose ForAug, a novel data augmentation scheme for image classification which makes the translation equivariance of CNNs explicit in the training data by recombining foreground objects at varying positions with different backgrounds.
-Applying ForAug to ImageNet gives rise to ForNet, a novel dataset that enables this data augmentation with with fine-grained control over the image composition.
-Recognizing that Transformers need to learn the spatial relationships from data, since they are not inherently translation invariant, and in general are usually trained on larger datasets [24], we separate the foreground objects in ImageNet from their backgrounds, using an open-world object detector [37], and fill in the background in a plausible way using an object removal model [43, 45].
-This allows us to recombine any foreground object with any background on the fly, creating a highly diverse training set.
-During recombination, we can control important parameters, like the size and position of the foreground object, to help the model learn the spatial invariances necessary for image classification.
-We show that training on ForNet instead of ImageNet increases the model accuracy of Transformers by up to 4.5 p.p.
-on ImageNet and an up to $39.3\%$ reduction in error rate on downstream tasks.
-Additionally, ForAug is a useful tool for analyzing model behavior and biases, when used during the evaluation phase.
-We utilize our control over the image distribution to quantify a model’s background robustness (by varying the choice of background), foreground focus (by leveraging our knowledge about the placement of the foreground object), center bias (by controlling the object’s position), and size bias (by controlling object size).
-These analyses provide insights into model behavior and biases, which is crucial for model deployment and future robustness optimizations.
-We show that training on ForNet, instead of ImageNet, significantly reduces all of these biases, completely removing the models’ dependence on the background distribution.
-We make our code for ForAug and the ForNet-dataset publicly available [^1] https://github.com/tobna/ForAug to facilitate further research.
+In particular, the self-attention mechanism, unlike a CNN, is not translation equivariant [38, 8], meaning that the model is not designed to understand the spatial relationships between pixels.
+Recognizing that Transformers need to learn spatial relationships directly from data, we propose ForAug, a data augmentation method that makes these relationships explicit by recombining foreground objects with diverse backgrounds.
+Thus, ForAug goes beyond existing image compositions and encodes desired invariances directly into the training data (see Figure 1).
+Applying ForAug to a dataset like ImageNet is a two-step process: (1) We separate the foreground objects in ImageNet from their backgrounds, using an open-world object detector [37] and fill in the background in a neutral way using an object removal model [43, 45].
+(2) This allows us to then recombine any foreground object with any background on the fly, creating a highly diverse training set.
+By exploiting the control over foreground size and position during recombination, ForAug explicitly teaches spatial invariances that image classification models typically must learn implicitly.
+We show that using ForAug additionally to strong traditional data augmentation increases the model accuracy of Transformers by up to 4.5 p.p.
+on ImageNet and reduces the error rate by up to $7.3$ p.p.
+in downstream tasks.
+Beyond training, ForAug becomes a diagnostic tool for analyzing model behavior and biases, when used during evaluation.
+We utilize our control over the image distribution to measure a model’s background robustness (by varying the choice of background), foreground focus (by leveraging our knowledge about the placement of the foreground object), center bias (by controlling position), and size bias (by controlling size).
+These analyses provide valuable insights into model behavior and biases, which is crucial for model deployment and future robustness optimizations.
+We show that training using ForAug significantly reduces all of these biases.
+We make our code for ForAug and the output of ForAug’s segmentation phase on ImageNet publicly available [^1] Link will go here.
+to facilitate further research.
 Contributions
-  • We propose ForAug, a novel data augmentation scheme, that recombines objects and backgrounds to train Transformers for image classification.
-  • We show that training on ForNet, the ImageNet instantiation of ForAug, leads to 4.5 p.p. improved accuracy on ImageNet and 7.3 p.p. on downstream tasks.
-  • We propose novel ForAug-based metrics to analyze and quantify fine-grained biases trained models: Background Robustness, Foreground Focus, Center Bias, and Size Bias.
-Training on ForNet, instead of ImageNet, significantly reduces these biases.
+  • We propose ForAug, a novel data augmentation scheme, that recombines objects and backgrounds.
+ForAug allows us to move beyond the (possibly biased) image compositions in the dataset while preserving label integrity.
+  • We show that training a standard ViT using ForAug leads to up to 4.5 p.p. improved accuracy on ImageNet-1k and 7.3 p.p. on downstream tasks.
+  • We propose novel ForAug-based metrics to analyze and quantify fine-grained biases of trained models: Background Robustness, Foreground Focus, Center Bias, and Size Bias.
+We show that ForAug significantly reduces these biases by encoding invariance that benefits ViT into the training data.
 2 Related Work
@@
 Traditional augmentation strategies rely on simple geometric or color-space transformations like cropping, flipping, roatation, blurring, color jittering, or random erasing [64] to increase the diversity of the training data without changing their semantic meaning.
-With the advent of Transformers, new data augmentation operations like PatchDropout [30] have been proposed.
+With the advent of Vision Transformers, new data augmentation operations like PatchDropout [29] have been proposed.
 Other transformations like Mixup [63], CutMix [61], or random cropping and patching [46] combine multiple input images.
-These simple transformations are usually bundled to form more complex augmentation policies like AutoAugment [5] and RandAugment [6], which automatically search for optimal augmentation policies or 3-augment [48] which is optimized to train a ViT.
-For a general overview of data augmentation techniques for image classification, we refer to [42, 56].
-We build upon these general augmentation techniques by introducing a novel approach to explicitly separate and recombine foregrounds and backgrounds for image classification.
-Our approach is used in tandem with traditional data augmentation techniques to improve model performance and reduce biases.
+These simple transformations are usually bundled to form more complex augmentation policies like AutoAugment [4] and RandAugment [5], or 3-augment [48] which is optimized to train a ViT.
+For a general overview of data augmentation techniques for image classification, we refer to Shorten and Khoshgoftaar [42], Xu et al.
+[58].
+We build upon these general augmentations by introducing a novel approach to explicitly separate objects and backgrounds for image classification, allowing us to – unlike these basic transformations – move beyond dataset image compositions.
+Our approach is used additionally to strong traditional techniques to improve performance and reduce biases.
 Copy-Paste Augmentation
-The copy-paste augmentation [14], which is used for object detection [41, 14] and instance segmentation [53, 28], involves copying segmented objects from one image and pasting them onto another.
-While typically human-annotated segmentation masks are used to extract the foreground objects, other foregound sources have been explored, like 3D models [19] and pretrained object-detection models for use on objects on white background [11] or synthetic images [12].
-DeePaste [53] focuses on using inpainting for a more seamless integration of the pasted object.
-Unlike these methods, ForNet focuses on image classification.
-While for detection and segmentation, objects are pasted onto another image (with a different foreground) or on available or rendered background images of the target scene, we extract foreground objects and fill in the resulting holes in the background in a semantically neutral way.
-This way, we can recombine any foreground object with a large variety of neutral backgrounds from natural images, enabling a controlled and diverse manipulation of image composition.
+The copy-paste augmentation [13], which is used only for object detection [41, 13] and instance segmentation [54, 27], involves copying segmented objects from one image and pasting them onto another.
+While typically human annotated segmentation masks are used to extract the foreground objects, other foregound sources have been explored, like 3D models [18] and pretrained object-detection models for use on objects on white background [10] or synthetic images [11].
+[20] apply copy-paste as an alternative to CutMix in image classification, but they do not shift the size or position of the foregrounds and use normal dataset images as backgrounds.
+Unlike prior copy-paste methods that overlay objects, ForAug extracts foregrounds and replaces their backgrounds with semantically neutral fills, thereby preserving label integrity while enabling controlled and diverse recombination.
+Figure 2: Overview of ForAug.
+The data creation consists of two stages: Segmentation (offline, Section 3.1), where we segment the foreground objects from the background and fill in the background.
+Recombination (online, Section 3.2), where we combine the foreground objects with different backgrounds to create new samples.
+After recombination, we apply strong, commonly used augmentation policies.
 Model robustness evaluation
@@
 Stylized ImageNet [12] investigates the impact of texture changes.
-ImageNet-9 [55] explores background variations using segmented images, but the backgrounds are often artificial.
-In contrast to these existing datasets, which are used only for evaluation, ForNet provides fine-grained control over foreground object placement, size, and background selection, enabling a precise and comprehensive analysis of specific model biases within the context of a large-scale, real-world image distribution.
-As ForNet also provides controllable training set generation, it goes beyond simply measuring robustness to actively improving it through training.
-3 RecombiNet (Method)
-Figure 2: Overview of ForNet.
-The data creation consists of two stages: (1, offline) Segmentation, where we segment the foreground objects from the background and fill in the background.
-(2, online) Recombination, where we combine the foreground objects with different backgrounds to create new samples.
-We introduce ForAug, a data augmentation scheme designed to enhance Transformer training by explicitly separating and recombining foreground objects and backgrounds.
-ForAug involves two stages: Segmentation and Recombination, both visualized in Figure 2.
-Segmentation
+ImageNet-9 [57] explores background variations using segmented images, but backgrounds are often artificial.
+In contrast to these existing datasets, which are used only for evaluation, ForAug provides fine-grained control over foreground object placement, size, and background selection, enabling a precise and comprehensive analysis of specific model biases within the context of a large-scale, real-world image distribution.
+As ForAug also provides controllable training set generation, it goes beyond simply measuring robustness to actively improving it through training.
+3 ForAug (Method)
+We introduce ForAug, a data augmentation designed to enhance Transformer training by embedding spatial invariances–which Transformers would otherwise need to learn implicitly–directly into the training data.
+ForAug comprises two distinct stages: Segmentation and Recombination.
+Both stages are illustrated in Figure 2.
+3.1 Segmentation
 The segmentation stage isolates the foreground objects and their corresponding backgrounds.
-We then fill in the background in a visually plausible way [43] using a pretrained object-removal model.
+We then fill the background using a pretrained object-removal model, producing visually plausible [43], neutral scenes ready for recombination.
 This stage is computed once offline and the results are stored for the recombination stage.
@@
 We derive the <object category> from the WordNet hierarchy, using the immediate hypernym.
-We iteratively extract up to $n$ foreground masks for each dataset-image, using different more and more general prompts based on the more general synsets of WordNet (e.g.
+We iteratively extract $n$ foreground masks for each dataset-image, creating prompts by going one hypernym up the WordNet-tree each step (e.g.
 “a sorrel, a type of horse”, “a horse, a type of equine”, …).
@@
 We select the best mask per image (according to Equation 1) in a later filtering step, described below.
-An inpainting model that is specifically optimized to remove objects from images, such as LaMa [45] or Attentive Eraser [43], is used to inpaint the foreground regions in the backgrounds.
-To ensure the quality of the foreground and background images (for each dataset-image), we select a foreground/background pair from the $\leq n$ variants we have extracted and infilled in the previous steps.
-Using an ensemble of six ViT, ResNet, and Swin Transformer models pretrained on the original dataset, we select the foreground/background pair that maximizes foreground performance while minimizing the performance on the background and size of the foreground according to:
+First, an inpainting model that is specifically optimized to remove objects from images, such as LaMa [45] or Attentive Eraser [43], is used to inpaint the foreground regions in the backgrounds.
+Then, to ensure the quality of the foregrounds and the neutral background images, we select a foreground/background pair (for each dataset-image) from the $\leq n$ variants we have extracted and infilled in the previous steps.
+Using an ensemble $E$ of six ViT, ResNet, and Swin Transformer models pretrained on the original dataset, we select the foreground/background pair that maximizes foreground performance while minimizing the performance on the background and size of the foreground.
+For each model $m\in E$, we predict the score of the ground truth class $c$ on the foreground $\mathrm{fg}$ and background $\mathrm{bg}$ and weigh these with the size $\operatorname{size}(\cdot)$ in number of pixels according to:
 $\begin{split}\text{score}(\mathrm{fg},\mathrm{bg},c)&=\log\left(\frac{1}{\left|E\right|}\sum_{m\in E}\mathbb{P}[m(\mathrm{fg})=c]\right)\\ &+\log\left(1-\frac{1}{\left|E\right|}\sum_{m\in E}\mathbb{P}[m(\mathrm{bg})=c]\right)\\ &+\lambda\log\left(1-\left|\frac{\operatorname{size}(\mathrm{fg})}{\operatorname{size}(\mathrm{bg})}-\varepsilon\right|\right).\end{split}$ (1)
-Here, $E$ is the ensemble of models and $m$ is a pretrained model, $c$ is the correct foreground class, $\mathrm{fg}$, and $\mathrm{bg}$ are the foreground and background and $\operatorname{size}(\cdot)$ is the size in number of pixels.
-We ran a hyperparameter search using a manually annotated subset of foreground/background variants to find the factors in Equation 1: $\lambda=2$ and $\varepsilon=0.1$.
-The optimal foreground size of $10\%$ of the full image balances the smallest possible foreground size that encompasses all the respective class information in the image with still conveying the foreground information after pasting it onto another background.
-This filtering step ensures we segment all the relevant foreground objects.
-Finally, we filter out backgrounds that are more than $80\%$ infilled, as these tend to be overly synthetic, plain and don’t carry much information (see Appendix C).
-We ablate this choice in Section 4.1.
+We run a hyperparameter search using a manually annotated subset of foreground/background variants to find the factors in Equation 1: $\lambda=2$ and $\varepsilon=0.1$.
+Finally, we filter out backgrounds that are largely infilled, as these tend to be overly synthetic and do not carry much information (see the supplementary material).
+Although the segmentation stage is computational overhead, it is a one-time cost with results that can be reused across experiments (see the supplementary material for details).
 In summary, we factorize the dataset into a set of foreground objects with a transparent background and a set of diverse backgrounds per class.
-The next step is to recombine them as data augmentation before applying common data augmentation operations during training.
-Recombination
-The recombination stage, which is performed online, combines the foreground objects with different backgrounds to create new training samples.
-For each object, we follow the pipeline of: Pick an appropriate background, resize it to a fitting size, place it in the background image, smooth the transition edge, and apply other data augmentations.
+The next step is to recombine these, before applying other common data augmentation operations during training.
+3.2 Recombination
+The recombination stage, performed online during training, combines the foreground objects with different backgrounds to create new training samples.
+For each object, we follow the pipeline of: Pick an appropriate background, resize it to a fitting size, and place it in the background image.
+Through this step, we expose the model to variations beyond the image compositions of the dataset.
 For each foreground object, we sample a background using one of the following strategies: (1) the original image background, (2) the set of backgrounds from the same class, or (3) the set of all possible backgrounds.
@@
 The selected foreground is resized based on its relative size within its original image and the relative size of the original foreground in the selected background image.
-The final size is randomly selected from a 30% range around upper and lower limits ($s_{u}$ and $s_{l}$), based on the original sizes:
-$s\sim\mathcal{U}\left[(1-0.3)s_{l},(1+0.3)s_{u}\right].$ (2)
+The final size is randomly selected from a 30% range around upper and lower limits ($s_{u}$ and $s_{l}$), based on the original sizes.
 To balance the size of the foreground and that of the backgrounds original foreground, the upper and lower limit $s_{u}$ and $s_{l}$ are set to the mean or range of both sizes, depending on the foreground size strategy: mean or range.
 The resized foreground is then placed at a random position within the background image.
-This position is sampled from a generalization of the Bates distribution [2] with parameter $\eta\in\mathbb{N}$, visualized in Figure 3.
-We choose the bates distribution, as it presents an easy way to sample from a bounded domain with just one hyperparameter that controls the concentration of the distribution.
-$\eta=1$ corresponds to the uniform distribution; $\eta>1$ concentrates the distribution around the center; and for $\eta<-1$, the distribution is concentrated at the borders.
 To more seamlessly integrate the foreground, we apply a Gaussian blur with $\sigma\in[\frac{\sigma_{\text{max}}}{10},\sigma_{\text{max}}]$}, inspired by the standard range for the Gaussian blur operation in [48], to the foreground’s alpha-mask.
 We can apply standard data augmentation techniques in two modes: Either we apply all augmentations to the recombined image, or we apply the cropping and resizing to the background only and then apply the other augmentations after recombination.
-The second mode ensures the foreground object remains fully visible, while the first mode mirrors standard data augmentation practices.
+The first mode mirrors standard augmentation practice, whereas the second one ensures the foreground object remains fully visible.
 We experiment with a constant mixing ratio, or a linear or cosine annealing schedule that increases the amount of images from the original dataset over time.
-The mixing ratio acts as a probability of selecting an image from the original dataset; otherwise, an image with the same foreground is recombined using ForAug.
-Thus, we still ensure each foreground is seen once per epoch.
+The mixing ratio acts as a probability of selecting an image from the original dataset; otherwise, an image with the same foreground is recombined using ForAug, ensuring each object is seen once per epoch.
+The recombination stage is designed to be parallelized on the CPU during training and thus does not impact training time (see supplementary material for details).
 4 Experiments
-We conduct a comprehensive suit of experiments to validate the effectiveness of our approach.
-We compare training on ForNet, the ImageNet instantiation of ForAug, to training on ImageNet for 7 different models.
-Furthermore, we assess the impact of using ForNet for pretraining on multiple fine-grained downstream datasets.
-Additionally, we use ForAug’s control over the image distribution to quantify some model behaviors and biases.
+We conduct a comprehensive suit of experiments to validate the effectiveness of our approach, comparing ImageNet-training with and without ForAug for 10 different models.
+Furthermore, we assess the impact of using ForAug for pretraining on multiple fine-grained downstream datasets.
+Finally, we exploit ForAug’s control over the image distribution to quantify model behaviors and biases.
+We always report the mean and standard deviation of three independent training runs.
 4.1 Design Choices of ForAug
-We start by ablating the design choices of ForAug.
-For this, we revert to TinyImageNet [26], a subset of ImageNet containing 200 categories with 500 images each, and TinyForNet, a version of ForAug derived from TinyImageNet.
-Table 1 presents the results of these ablations.
-Dataset Detect. Infill FG. Augmentation BG. BG. edge original TinyImageNet Accuracy
-prompt Model size Order strategy pruning smoothing image mixing ViT-Ti [%] ViT-S [%]
+We start by ablating the design choices of ForAug on TinyImageNet [25], a subset of ImageNet containing 200 categories with 500 images each.
+Table 1 presents ablations for segmentation and Table 2 for recombination.
+Table 1: Ablation of the design decisions in the segmentation phase of ForAug on TinyImageNet.
+The first line is our baseline, while the other lines are using ForAug.
+We use basic settings with the same background strategy during recombination for this experiment.
+Detect. Prompt Infill Model TinyImageNet Accuracy [%]
+ViT-Ti ViT-S
+TinyImageNet $66.1\pm 0.5$ $68.3\pm 0.7$
+specific LaMa [45] $65.5\pm 0.4$ $71.2\pm 0.5$
+general LaMa [45] $66.4\pm 0.6$ $72.9\pm 0.6$
+general Att. Eraser [43] $67.5\pm 1.2$ $72.4\pm 0.5$
+Table 2: Ablation of the recombination phase of ForAug on TinyImageNet (top) and ImageNet (bottom). The first experiments use the initial segmentation settings with LaMa [45].
+FG.
+size
+Augment.
+Order
+BG
+Strat.
+BG.
+Prune
+Original
+Mixing
+Edge
+Smooth.
+Accuracy [%]
+ViT-Ti ViT-S
 TinyImageNet $66.1\pm 0.5$ $68.3\pm 0.7$
@@
 […]
+ImageNet - $79.1\pm 0.1$
 […]
@@
 […]
-[…]
-[…]
-[…]
-Table 1: Ablation of design decisions of TinyForNet on TinyImageNet and ForNet on ImageNet.
 Prompt.
@@
 Here, the general prompt, which contains the class and the more general object category, outperforms only having the class name (specific).
-Inpainting. Attentive Eraser [43] produces superior results compared to LaMa [45] (see Appendix B for examples).
+Inpainting. Among inpainting models, Attentive Eraser [43] produces slightly better results compared to LaMa [45] (+0.5+0.5 p.p. on average). For inpainting examples, see the supplementary material.
 Foreground size significantly impacts performance.
@@
 Background pruning.
-When it comes to the choice of backgrounds to use, we test two pruning thresholds ($t_{\text{prune}}$) to exclude backgrounds with excessive inpainting.
+When it comes to the backgrounds to use, we test different pruning thresholds ($t_{\text{prune}}$) to exclude backgrounds with large inpainting.
 A threshold of $t_{\text{prune}}=1.0$ means that we use all backgrounds that are not fully infilled.
 Varying $t_{\text{prune}}$ has minimal impact.
-Therefore, we choose $t_{\text{prune}}=0.8$ to exclude predominantly artificial backgrounds.
-Similarly, applying edge smoothing to foreground masks with Gaussian blurring actually hurts performance on TinyForNet, but slightly improves it on ForNet.
-Mixing ForNet with the original ImageNet data proves crucial.
-While constant and linear mixing schedules improve performance over no mixing by $2-3$ p.p.
-compared to only using TinyForNet, the cosine annealing schedule yields the best results, boosting accuracy by another $0.5-1$ p.p.
+We choose $t_{\text{prune}}=0.8$ to exclude predominantly artificial backgrounds.
+Mixing ForAug-augmented samples with the original ImageNet data proves crucial.
+While constant and linear mixing schedules improve performance over no mixing by 232-3 p.p.
+compared to only augmented samples, the cosine annealing schedule proves optimal, boosting accuracy by 343-4 $p.$
+Edge smoothing.
+We evaluate the impact of using Gaussian blurring to smooth the edges of the foreground masks.
+For larger models, this gives us a slight performance boost on the full ImageNet (second to last line in Table 2).
 Background strategy.
@@
 For ViT-Ti, the latter two strategies perform comparably, while ViT-S benefits from the added diversity of using any background.
-The same is true when training on the full (ImageNet) version of ForNet.
-Figure 3: Plot of the probability distribution function (PDF) of the extended Bates distribution for different parameters $\eta$.
-Higher values of $\eta$ concentrate the distribution around the center.
-Training Set/ Bates Parameter TIN TinyForNet
-$\eta=-3$ $-2$ $/-1$ 1 $2$ $3$
-TinyImageNet 68.9 60.5 60.2 60.8 62.6 63.1
+The same is true when training on the full ImageNet.
+Table 3: Accuracy of ViT-S on TinyImageNet (TIN) in percent using ForAug with different foreground position distributions by varying the Bates parameter $\eta$.
+The best performance is achieved when using the uniform distribution ($\eta=1$) for training.
+Bates Parameter during training TIN w/o ForAug TIN w/ ForAug
+$\eta=-3$ 2-2 1/11/-1 22 33
+Baseline 68.9 60.5 60.2 60.8 62.6 63.1
 $\eta=-3$ 71.3 79.3 79.5 79.1 79.3 79.1
@@
 $\eta=3$ 71.4 77.2 76.9 78.6 79.6 79.7
-Table 2: Accuracy of ViT-S trained on TinyImageNet (TIN) and TinyForNet with different foreground position distributions by varying the parameter of a Bates distribution $\eta$.
-The best performance is achieved using the uniform distribution ($\eta=1$).
 Foreground position.
-Finally, we analyze the foreground object’s positioning in the image.
-We utilize an extended Bates distribution to sample the position of the foreground object.
-The Bates distribution [2] with parameter $\eta\geq 1$ is the mean of $\eta$ independent uniformly distributed random variables [20].
-Therefore, the larger $\eta$, the more concentrated the distribution is around the center.
-We extend this concept to $\eta\leq-1$ by defining ${X\sim\text{Bates}(\eta):\Leftrightarrow s(X)\sim\text{Bates}(-\eta)}$ for $\eta\leq 1$ with $s$ being the sawtooth function on $[0,1]$:
-$s(x)=\begin{cases}x+0.5&\text{if }0<x<0.5\\ x-0.5&\text{if }0.5\leq x\leq 1\end{cases}$ (3)
-Note that $s\circ s=\mathrm{id}$ on $[0,1]$.
-This way, distributions with $\eta\leq-1$ are more concentrated around the borders.
-$\eta=1$ and $\eta=-1$ both correspond to the uniform distribution.
-The PDF of this extended Bates distribution is visualized in Figure 3.
-When sampling more towards the center of the image, the difficulty of the task is reduced, which then reduces the performance on TinyImageNet.
-This is reflected in the performance when evaluating on TinyForNet with $\eta=2$ and $\eta=3$ compared to $\eta=-1/1$.
+Finally, we analyze the foreground object’s positioning in the image, using a generalization of the Bates distribution [1] with parameter $\eta\in\mathbb{Z}$.
+The Bates distribution presents an easy way to sample from a bounded domain with just one hyperparameter that controls its concentration.
+$\eta=1/-1$ corresponds to the uniform distribution; $\eta>1$ concentrates the distribution around the center; and for $\eta<-1$, the distribution is concentrated at the borders (see supplementary material for details).
+When sampling more towards the center of the image, the difficulty of the task is reduced, which reduces performance on TinyImageNet (Table 3).
+This is reflected in the performance when evaluating using ForAug with $\eta=2$ and $\eta=3$ compared to $\eta=-1/1$.
 We observe a similar reduction for $\eta<-1$.
-This experiment is conducted using the LaMa infill model.
-Dataset Classes Training Images Validation Images
-TinyImageNet 200 100,000 10,000
-TinyForNet 200 99,404 9,915
-ImageNet 1,000 1,281,167 50,000
-ForNet 1,000 1,274,557 49,751
-Table 3: Dataset statistics for TinyImageNet, TinyForNet, ImageNet, and ForNet. For ForNet and TinyForNet we report the number of foreground/background pairs.
-After fixing the optimal design parameters in Table 1 (last row), we construct the full ForNet dataset using the entire ImageNet dataset.
-Table 3 compares the dataset statistics of ImageNet and ForNet.
-The slightly reduced image count in ForNet is due to instances where Grounded SAM failed to produce valid object detections.
+Table 4: Dataset statistics for TinyImageNet and ImageNet with and without ForAug. For ForAug we report the number of foreground/background pairs.
+Dataset Classes
+Training
+Images
+Validation
+Images
+TinyImageNet $200$ 100 000100\,000 10 00010\,000
+TinyImageNet + ForAug $200$ 99 40499\,404 $9915$
+ImageNet $1000$ 1 281 1671\,281\,167 50 00050\,000
+ImageNet + ForAug $1000$ 1 274 5571\,274\,557 49 75149\,751
+After fixing the optimal design parameters in Tables 1 and 2 (last rows), we run ForAug’s segmentation step on the entire ImageNet dataset.
+Table 4 shows the resulting dataset statistics.
+The slightly reduced image count for ForAug is due to instances where Grounded SAM fails to produce valid segmentation masks.
 4.2 Image Classification Results
-Model ImageNet Accuracy when trained on Delta
-ImageNet ForNet
-ViT-S $79.1\pm 0.1$ $81.4\pm 0.1$ +2.3
-ViT-B $77.6\pm 0.2$ $81.1\pm 0.4$ +3.5
-ViT-L $75.3\pm 0.4$ $79.8\pm 0.1$ +4.5
-Swin-Ti $77.9\pm 0.2$ $79.7\pm 0.1$ +1.8
-Swin-S $79.4\pm 0.1$ $80.6\pm 0.1$ +1.2
-ResNet-50 $78.3\pm 0.1$ $78.8\pm 0.1$ +0.5
-ResNet-101 $79.4\pm 0.1$ $80.4\pm 0.1$ +1.0
-Table 4: ImageNet results of models trained on ForNet and on ImageNet directly. ForNet improves the performance of all models in our test.
-Table 4 compares the ImageNet performance of models trained on ForNet and ones trained directly on ImageNet.
-We adopt the training setup of [33] and [48] (details in Appendix A) for training ViT [10], Swin [31] and ResNet [16] models.
-Notably, ForNet improves performance across all tested architectures, including the ResNet models (up to $1$ p.p.), demonstrating benefits beyond Transformers.
-For Transformer models, we observe improvements from $1.2$ p.p.
+Table 5: ImageNet results of models trained on ImageNet with and without ForAug. ForAug improves the performance of most models, with a larger gain for larger models.
+Model ImageNet Accuracy [%] Delta
+w/o ForAug w/ ForAug
+ViT-S $79.1\pm 0.1$ $81.4\pm 0.1$ +2.3+2.3
+ViT-B $77.6\pm 0.2$ $81.1\pm 0.4$ +3.5+3.5
+ViT-L $75.3\pm 0.4$ $79.8\pm 0.1$ +4.5+4.5
+DeiT-S $80.1\pm 0.1$ $80.0\pm 0.3$ 0.1-0.1
+DeiT-B $81.9\pm 0.3$ $81.9\pm 0.2$ $\pm 0.0$
+DeiT-L $79.3\pm 2.3$ $82.4\pm 0.1$ +3.1+3.1
+Swin-Ti $77.9\pm 0.2$ $79.7\pm 0.1$ +1.8+1.8
+Swin-S $79.4\pm 0.1$ $80.6\pm 0.1$ +1.2+1.2
+ResNet-50 $78.3\pm 0.1$ $78.8\pm 0.1$ +0.5+0.5
+ResNet-101 $79.4\pm 0.1$ $80.4\pm 0.1$ +1.0+1.0
+Table 5 compares the ImageNet performance of models trained with and without ForAug.
+We adopt the training setup of [32] and [48] for training ViT [9], Swin [30] and ResNet [15] (representing CNNs) models as well as the setup of DeiT [47] for that model.
+Both setups are using strong data augmentations like RandAugment, CutMix, and Mixup optimized for Transformers (details in supplementary material).
+Notably, ForAug improves performance across all tested architectures, including the ResNet models, demonstrating benefits beyond Transformers.
+For DeiT we only observe benefits on ImageNet for the larger models.
+For other transformers, we observe improvements from $1.2$ p.p.
 to $4.5$ p.p.
-This improvement is more substantial for the larger models, with ViT-L gaining $4.5$ p.p.
-in accuracy.
-ForNet’s improvements mostly counteract the drop in performance due to overfitting for large models.
-When training on ImageNet, this drop is $3.8$ p.p.
-from ViT-S to ViT-L, while for ForNet it is reduced to $1.6$ p.p.
-Model Aircraft Cars Flowers Food Pets
-ViT-S @ ImageNet $72.4\pm 1.0$ $89.8\pm 0.3$ $94.5\pm 0.2$ $89.1\pm 0.1$ $93.8\pm 0.2$
-ViT-S @ ForNet $78.6\pm 0.5$ $92.2\pm 0.2$ $95.5\pm 0.2$ $89.6\pm 0.1$ $94.5\pm 0.2$
-+6.2 +2.4 +1.0 +0.5 +0.7
-ViT-B @ ImageNet $71.7\pm 0.5$ $90.0\pm 0.2$ $94.8\pm 0.4$ $89.8\pm 0.2$ $94.1\pm 0.4$
-ViT-B @ ForNet $79.0\pm 2.2$ $93.3\pm 0.1$ $96.5\pm 0.1$ $90.9\pm 0.1$ $95.1\pm 0.4$
-+7.3 +3.3 +1.7 +1.1 +1.0
-ViT-L @ ImageNet $72.1\pm 1.0$ $88.8\pm 0.3$ $94.4\pm 0.3$ $90.1\pm 0.2$ $94.2\pm 0.4$
-ViT-L @ ForNet $77.6\pm 1.2$ $89.1\pm 0.2$ $96.6\pm 0.1$ $91.3\pm 0.1$ $95.1\pm 0.1$
-+5.5 +0.3 +2.2 +1.2 +0.9
-Swin-Ti @ ImageNet $77.0\pm 0.1$ $91.3\pm 0.6$ $95.9\pm 0.1$ $90.0\pm 0.2$ $94.2\pm 0.1$
-Swin-Ti @ ForNet $81.1\pm 0.8$ $92.8\pm 0.4$ $96.2\pm 0.1$ $90.4\pm 0.3$ $94.8\pm 0.5$
-+4.1 +2.5 +0.3 +0.4 +0.6
-Swin-S @ ImageNet $75.7\pm 1.4$ $91.0\pm 0.3$ $95.9\pm 0.5$ $91.1\pm 0.2$ $94.4\pm 0.1$
-Swin-S @ ForNet $81.4\pm 0.2$ $93.1\pm 0.2$ $96.3\pm 0.3$ $91.2\pm 0.2$ $94.9\pm 0.3$
-+5.7 +2.1 +1.4 +0.1 +0.5
-ResNet-50 @ ImageNet $78.2\pm 0.5$ $89.8\pm 0.2$ $91.7\pm 0.4$ $84.4\pm 0.2$ $93.7\pm 0.3$
-ResNet-50 @ ForNet $80.3\pm 0.4$ $90.4\pm 0.2$ $91.7\pm 0.2$ $84.5\pm 0.2$ $93.7\pm 0.3$
-+2.1 +0.6 $\pm$0 +0.1 $\pm$0
-ResNet-101 @ ImageNet $78.4\pm 0.6$ $90.3\pm 0.1$ $91.2\pm 0.5$ $86.0\pm 0.2$ $94.3\pm 0.2$
-ResNet-101 @ ForNet $81.4\pm 0.5$ $91.3\pm 0.1$ $92.9\pm 0.2$ $86.3\pm 0.1$ $94.0\pm 0.3$
-+3.0 +1.3 +1.7 +0.3 -0.3
+with increasing gains for larger models.
+ForAug’s improvements counteract the drop in performance for increasing model sizes.
+Without ForAug this drop is $3.8$ p.p.
+(ViT-S to L), while with ForAug it is reduced to $1.6$ p.p.
+For DeiT there is a drop of $0.8$ p.p.
+from small to large while when using ForAug there is a gain of $2.4$ $p.$
+Table 6: Comparison of ForAug and simple Copy-Paste methods. We train ViT-S on ImageNet using the same 3-augment data augmentation on top of the copy-paste augmentation.
+Augmentation labels
+Accuracy [%]
+Delta
+to Prev.
+Baseline + Simple Copy-Paste bg $31.3\pm 0.6$
++ mixed labels fg + bg $32.0\pm 0.8$ +0.7+0.7
++ fg labels fg $31.6\pm 0.9$ 0.4-0.4
++ range foreground size variation fg $43.0\pm 1.2$ +11.4+11.4
++ infilled backgrounds fg $68.7\pm 0.2$ +25.7+25.7
++ cos mixing strategy fg $81.2\pm 0.1$ +12.5+12.5
++ edge smoothing fg $81.3\pm 0.1$ +0.1+0.1
++ background pruning== ForAug fg $81.4\pm 0.1$ +0.1+0.1
+Comparison to Simple Copy-Paste.
+We compare ForAug to a simple adaption of the Copy-Paste augmentation inspired by [11, 13, 41] in Table 6.
+Contrary to semantic segmentation we do not have foreground masks available.
+Thus, we paste the extracted foreground objects from ForAug’s segmentation stage onto normal ImageNet images.
+We observe 3 large jumps in accuracy: (1) From our range foreground size variation (+11.4%), (2) from using our infilled backgrounds instead of images from the dataset (+25.7%), and (3) from our cos mixing strategy with non-augmented images (+12.5%).
+ForAug’s changes to the naive copy-paste augmentation are thus imperative for good classification performance.
 Table 7: Downstream accuracy in percent when finetuning on other datasets.
-Models were pretrained on ForNet and ImageNet.
-Pretraining on ForNet increases Transformer downstream accuracy on all datasets.
-To assess the transferability of ForNet-trained models, we finetune models pretrained on ImageNet and ForNet on five fine-grained datasets: FGVC-Aircraft [32], Stanford Cars [7], Oxford Flowers [34], Food-101 [21], and Oxford-IIIT Pets [35].
-While for ResNets, the performance of both training datasets is about the same, for every Transformer, we see the accuracy improve on all downstream dataset by up to 7.3 p.p.
-and a reduction of error rate of up to $39.3\%$.
-In summary, these results demonstrate that the improved representation learning achieved by training on ForNet translates to superior performance not only on ImageNet, but also on a variety of fine-grained image classification tasks.
-4.3 Further Model Evaluation
-Beyond its use for training, ForNet’s unique properties and controlled data generation capabilities make it a powerful tool for analyzing model behavior and biases.
-Background Robustness
-Model Background Robustness when trained on Delta
-ImageNet ForNet
-ViT-S $0.73\pm 0.01$ $0.99\pm 0.01$ +0.26
-ViT-B $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
-ViT-L $0.70\pm 0.01$ $1.00\pm 0.01$ +0.30
-Swin-Ti $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
-Swin-S $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
-ResNet-50 $0.79\pm 0.01$ $0.99\pm 0.01$ +0.20
-ResNet-101 $0.79\pm 0.01$ $1.00\pm 0.01$ +0.21
-Table 6: Evaluation of the background robustness of models trained on ForNet and on ImageNet directly.
-Training on ForNet improves the background robustness of all model to $\approx 1.00$, meaning the model is indifferent to the choice of background.
+Models are pretrained on ImageNet with and without ForAug.
+Pretraining using ForAug increases transformer downstream accuracy.
+Model ForAug Aircraft Cars Flowers Food Pets
+ViT-S $72.4\pm 1.0$ $89.8\pm 0.3$ $94.5\pm 0.2$ $89.1\pm 0.1$ $93.8\pm 0.2$
+ViT-S $78.6\pm 0.5$ $92.2\pm 0.2$ $95.5\pm 0.2$ $89.6\pm 0.1$ $94.5\pm 0.2$
++6.2+6.2 +2.4+2.4 +1.0+1.0 +0.5+0.5 +0.7+0.7
+ViT-B $71.7\pm 0.5$ $90.0\pm 0.2$ $94.8\pm 0.4$ $89.8\pm 0.2$ $94.1\pm 0.4$
+ViT-B $79.0\pm 2.2$ $93.3\pm 0.1$ $96.5\pm 0.1$ $90.9\pm 0.1$ $95.1\pm 0.4$
++7.3+7.3 +3.3+3.3 +1.7+1.7 +1.1+1.1 +1.0+1.0
+ViT-L $72.1\pm 1.0$ $88.8\pm 0.3$ $94.4\pm 0.3$ $90.1\pm 0.2$ $94.2\pm 0.4$
+ViT-L $77.6\pm 1.2$ $89.1\pm 0.2$ $96.6\pm 0.1$ $91.3\pm 0.1$ $95.1\pm 0.1$
++5.5+5.5 +0.3+0.3 +2.2+2.2 +1.2+1.2 +0.9+0.9
+DeiT-S $75.3\pm 0.4$ $91.1\pm 0.2$ $94.8\pm 0.4$ $89.2\pm 0.2$ $92.4\pm 0.2$
+DeiT-S $76.8\pm 0.8$ $91.9\pm 0.2$ $95.2\pm 0.3$ $89.1\pm 0.2$ $92.3\pm 0.4$
++1.5+1.5 +0.8+0.8 +0.4+0.4 0.1-0.1 0.1-0.1
+DeiT-B $77.0\pm 1.2$ $92.9\pm 0.2$ $96.1\pm 0.2$ $91.2\pm 0.1$ $93.3\pm 0.4$
+DeiT-B $79.3\pm 0.3$ $93.1\pm 0.1$ $96.4\pm 0.2$ $91.3\pm 0.1$ $93.3\pm 0.1$
++2.3+2.3 +0.2+0.2 +0.3+0.3 +0.1+0.1 $\pm 0.0$
+DeiT-L $72.8\pm 5.5$ $92.8\pm 1.0$ $95.8\pm 1.5$ $90.5\pm 2.6$ $92.4\pm 2.0$
+DeiT-L $78.8\pm 0.8$ $93.8\pm 0.2$ $97.0\pm 0.2$ $92.0\pm 0.2$ $93.5\pm 0.2$
++6.0+6.0 +1.0+1.0 +1.2+1.2 +1.5+1.5 +1.1+1.1
+Swin-Ti $77.0\pm 0.1$ $91.3\pm 0.6$ $95.9\pm 0.1$ $90.0\pm 0.2$ $94.2\pm 0.1$
+Swin-Ti $81.1\pm 0.8$ $92.8\pm 0.4$ $96.2\pm 0.1$ $90.4\pm 0.3$ $94.8\pm 0.5$
++4.1+4.1 +2.5+2.5 +0.3+0.3 +0.4+0.4 +0.6+0.6
+Swin-S $75.7\pm 1.4$ $91.0\pm 0.3$ $95.9\pm 0.5$ $91.1\pm 0.2$ $94.4\pm 0.1$
+Swin-S $81.4\pm 0.2$ $93.1\pm 0.2$ $96.3\pm 0.3$ $91.2\pm 0.2$ $94.9\pm 0.3$
++5.7+5.7 +2.1+2.1 +1.4+1.4 +0.1+0.1 +0.5+0.5
+ResNet-50 $78.2\pm 0.5$ $89.8\pm 0.2$ $91.7\pm 0.4$ $84.4\pm 0.2$ $93.7\pm 0.3$
+ResNet-50 $80.3\pm 0.4$ $90.4\pm 0.2$ $91.7\pm 0.2$ $84.5\pm 0.2$ $93.7\pm 0.3$
++2.1+2.1 +0.6+0.6 $\pm 0.0$ +0.1+0.1 $\pm 0.0$
+ResNet-101 $78.4\pm 0.6$ $90.3\pm 0.1$ $91.2\pm 0.5$ $86.0\pm 0.2$ $94.3\pm 0.2$
+ResNet-101 $81.4\pm 0.5$ $91.3\pm 0.1$ $92.9\pm 0.2$ $86.3\pm 0.1$ $94.0\pm 0.3$
++3.0+3.0 +1.3+1.3 +1.7+1.7 +0.3+0.3 0.3-0.3
+Downstream tasks.
+To assess the transferability of ForAug-trained models, we finetune models pretrained on ImageNet with and without ForAug on five fine-grained datasets: FGVC-Aircraft [31], Stanford Cars [6], Oxford Flowers [33], Food-101 [21], and Oxford-IIIT Pets [34].
+In Table 7 we see transformer accuracies improve on all these datasets by up to 7.3 p.p.
+Notably, training with ForAug boosts the downstream performance of DeiT-S and DeiT-B, despite similar ImageNet results.
+This shows the improved representations from training with ForAug translate to gains beyond better ImageNet scores.
+4.3 Bias and Robustness Evaluation
+Beyond its use for training, ForAug’s unique properties and controlled data generation capabilities make it a powerful tool for analyzing behavior and biases of black-box models.
+Figure 3: Evaluation of background robustness on ImageNet + ForAug, ImageNet9 and CounterAnimal.
+We plot the in-distribution (top of arrow) and the out-of-distribution (bottom of arrow) accuracy when training with and without ForAug.
+We annotate each arrow with its length $\Delta$.
+Training with ForAug improves the background robustness of all transformers by mostly boosting the out-of-distribution accuracy.
+Background Robustness.
 We assess the robustness of models to shifts in the background distribution from a class-related background to any background.
-Background robustness is defined to be the ratio of accuracy on ForNet with same-class backgrounds to accuracy with any background:
-Background $\text{Background Robustness}=\frac{\text{Acc}(\textit{ForNet}_{\text{all}})}{\text{Acc}(\textit{ForNet}_{\text{same}})}$ (4)
-It represents the relative drop in performance under a background distribution shift.
-Table 6 presents the background robustness of various models.
-When trained on ImageNet, smaller models generally exhibit greater robustness to changes in the background distribution than larger models and ResNet is more robust than the tested Transformer models.
-Crucially, training on ForNet instead of ImageNet improves the background robustness of all models to $\approx 1.00$, meaning that these models are agnostic to the choice of background and only classify based on the foreground.
-These findings highlight the generalization benefits of ForNet.
-Foreground Focus
-Model Foreground Focus when trained on
-IN FN IN FN IN FN
-GradCam GradCam++ IG
-ViT-S $1.2\pm 0.1$ $2.3\pm 0.3$ $1.2\pm 0.1$ $2.1\pm 0.4$ $1.9\pm 0.1$ $2.7\pm 0.1$
-ViT-B $1.2\pm 0.1$ $2.4\pm 0.7$ $1.1\pm 0.1$ $2.1\pm 0.1$ $1.7\pm 0.1$ $2.7\pm 0.1$
-ViT-L $1.3\pm 0.1$ $1.6\pm 0.1$ $1.1\pm 0.1$ $1.3\pm 0.1$ $1.3\pm 0.1$ $2.6\pm 0.1$
-Swin-Ti $0.9\pm 0.1$ $0.7\pm 0.1$ $1.0\pm 0.3$ $0.7\pm 0.3$ $2.5\pm 01$ $4.8\pm 0.3$
-Swin-S $0.8\pm 0.1$ $0.7\pm 0.1$ $0.7\pm 0.1$ $0.7\pm 0.4$ $2.4\pm 0.1$ $4.6\pm 0.3$
-ResNet-50 $2.2\pm 0.1$ $2.7\pm 0.1$ $2.0\pm 0.1$ $2.9\pm 0.1$ $3.2\pm 0.1$ $4.9\pm 0.2$
-ResNet-101 $2.3\pm 0.1$ $2.8\pm 0.1$ $2.2\pm 0.1$ $3.0\pm 0.1$ $3.2\pm 0.1$ $4.8\pm 0.1$
-Table 7: Evaluation of the foreground focus using GradCam, GradCam++ and IntegratedGradients of models trained on ForNet (FN) and on ImageNet (IN) directly.
-Training on ForNet improves the foreground focus of almost all models.
-Leveraging our inherent knowledge of the foreground masks when using ForNet, as well as common XAI techniques [40, 4, 44], we can evaluate a model’s focus on the foreground object.
-We can directly evaluate ImageNet trained models, but this technique can also be extended to other datasets without relying on manually annotated foreground-masks.
-To evaluate the foreground focus, we employ Grad-CAM [40], Grad-CAM++ [4] or IntegratedGradients (IG) [44] to compute the per-pixel importance of an image for the model’s prediction.
+Figure 3 presents the background robustness results for three datasets: ImageNet with ForAug (all backgrounds vs.
+backgrounds of same class), ImageNet9 [57] (random backgrounds vs.
+original backgrounds), and CounterAnimal [51] (counter vs.
+common background).
+The top triangle of each arrow represents the in-distribution backgrounds and the bottom triangle represents the out-of-distribution ones.
+We follow ImageNet9 and CounterAnimal and assess the background robustness in terms of the accuracy gap when evaluating a model on images of normal background distribution compared to out-of-distribution backgrounds (length of each arrow; $\Delta$).
+Crucially, ForAug improves the background robustness of all models and across datasets, reducing the background-gap by boosting the performance on the out-of-background-distribution samples more than the in-distribution ones.
+These findings highlight the generalization benefits of ForAug to unusual image compositions.
+Figure 4: Evaluation of the foreground focus (Equation 2) using GradCam, GradCam++ and IntegratedGradients (IG) of models trained on ImageNet.
+Training with ForAug improves the foreground focus of almost all models.
+Foreground Focus.
+Leveraging our inherent knowledge of the foreground masks when using ForAug, as well as common XAI techniques [40, 3, 44], we can evaluate a model’s focus on the foreground object.
+We can directly evaluate ImageNet-trained models, but this technique can also be extended to other datasets without relying on manually annotated foreground masks.
+To evaluate the foreground focus, we employ Grad-CAM [40], Grad-CAM++ [3] and IntegratedGradients (IG) [44] to compute the per-pixel importance of an image for the model’s prediction.
 The foreground focus is defined to be the ratio of the foreground’s relative importance to its relative size in the image:
-[…]
-The foreground focus of a model is its average foreground focus over all test images.
-Table 7 presents our findings.
-Training on ForNet significantly increasees the foreground focus of ViT and ResNet across all metrics used.
-For Swin, the foreground focus stagnates when measured using GradCam and GradCam++, but almost doubles when using IG.
-Center Bias
-Model Center Bias when trained on Delta
-ImageNet ForNet
+$\text{FG Focus}(\text{img})=\frac{\text{Area}(\text{img})\hskip 3.0pt\text{Importance}(\text{fg})}{\text{Area}(\text{fg})\hskip 3.0pt\text{Importance}(\text{img})}$ (2)
+If all pixels uniformly receive the same importance value, the foreground focus is one.
+The foreground focus of a model is its average focus over all test images.
+Figure 4 presents our findings.
+Using ForAug significantly increases the foreground focus of ViT, DeiT and ResNet across all XAI metrics.
+We hypothesize Swin’s below-uniform foreground focus with GradCam is due to its specific implementation.
+Table 8: Accuracy relative to the center accuracy of multiple instantiations of the models when the foreground objects is in different cells of a $3\times 3$ grid.
+We calculate center bias according to Equation 3.
+Using ForAug significantly reduces models’ center bias.
+Model
+Center Bias [%] when trained
+Delta
+w/o ForAug w/ ForAug
 ViT-S [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-$0.255\pm 0.008$ $0.220\pm 003$0.035
+$5.5\pm 0.8 22$$2.0\pm 0.3$ 3.5-3.5
 ViT-B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-$0.254\pm 0.004$ $0.190\pm 0.002$0.064
+$25.4\pm 0.4$ $19.0\pm 0.2$ 6.4-6.4
 ViT-L [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-$0.243\pm 0.011$ $0.117\pm 0.007$0.126
+$24.3\pm 1.1$ $11.7\pm 0.7$ 12.6-12.6
+DeiT-S [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+$20.4\pm 0.2$ $21.2\pm 0.1$ +0.8+0.8
+DeiT-B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+$19.0\pm 0.7$ $19.0\pm 0.2$ $\pm 0.0$
+DeiT-L [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+$21.2\pm 0.2$ $18.0\pm 0.2$ 3.2-3.2
 Swin-Ti [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-$0.250\pm 0.007$ $0.165\pm 0.002$0.085
+$25.0\pm 0.7$ $16.5\pm 0.2$ 8.5-8.5
 Swin-S [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-$0.232\pm 0.001$ $0.156\pm 002$0.076
+$23.2\pm 0.1$ $15.6\pm 0.2$ 7.6-7.6
 ResNet50 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-$0.263\pm 0.003$ $0.197\pm 0.003$0.066
+$26.3\pm 0.3$ $19.7\pm 0.3$ 6.6-6.6
 ResNet101 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-$0.230\pm 0.003$ $0.199\pm 002$0.031
-Table 8: Evaluation of the position bias.
-We plot the accuracy relative to the center accuracy of multiple instantiations of the models when the foreground objects is in different cells a $3\times 3$ grid.
-Training on ForNet significantly reduces a models center bias.
-With ForNet we have unique control over the position of the foreground object in the image.
-This lets us quantify the center bias of ImageNet- and ForNet-trained models.
-We divide the image into a $3\times 3$ grid and evaluate model accuracy when the foreground object is in each of the $9$ grid cells.
+$23.0\pm 0.3$ $19.9\pm 0.2$ 3.1-3.1
+Center Bias.
+With ForAug we have unique control over the position of the foreground object in the image.
+This lets us quantify the center bias of models trained with and without ForAug.
+We divide the image into a $3\times 3$ grid and evaluate model accuracy when the (scaled-down) foreground object is in each of the $9$ grid cells.
 Each cell’s accuracy is divided by the accuracy in the center cell for normalization, which gives us the relative performance drop when the foreground is in each part of the image.
 The center bias is calculated as one minus the average of the minimum performance of a corner cell and the minimum performance of a side cell:
-Center $\begin{split}&\text{Center Bias}=\\ &\hskip 7.0pt1-\frac{\min\limits_{a,b\in\{0,2\}}\text{Acc}(\text{cell}_{(a,b)})+\min\limits_{\begin{subarray}{c}a=1\text{ or }b=1\\ a\neq b\end{subarray}}\text{Acc}(\text{cell}_{(a,b)})}{2\text{Acc}(\text{cell}_{(1,1)})}\end{split}$ (6)
+$\text{Center Bias}=1-\frac{\min\limits_{c\in\text{sides}}\text{Acc}(c)+\min\limits_{c\in\text{corners}}\text{Acc}(c)}{2\text{Acc}(c_{\text{center}})}$ (3)
 Table 8 visualizes the center bias of three instantiations of each model.
-Performance is generally highest in the center and the center top and bottom and center left and right cells, and lowest in the four corners.
+Performance is generally highest in the center and lowest in the four corners.
 Interestingly, ImageNet-trained models perform slightly better when the foreground object is on the right side of the image, compared to the left side, despite our use of random flipping with a probability of $0.5$ during training.
-Training on ForNet significantly reduces center bias across all models.
-This demonstrates that ForNet promotes a more uniform spatial attention distribution.
-Their accuracy is higher in the center left and right cells than in the center top and bottom ones, which is not the case for ImageNet-trained models.
-Size Bias
-Figure 4: Evaluation of the size bias of models trained on ForNet. We plot the accuracy relative to the accuracy when using the mean foreground size.
-Finally, we evaluate the impact of different-sized foreground objects on the accuracy.
+Using ForAug significantly reduces center bias across models, with a more uniform performance especially across the middle row.
+Thus, ForAug makes the model recognize objects across a wider spatial distribution, counteracting the center-bias of ImageNet.
+Figure 5: Evaluation of the size bias of models trained on ImageNet. We plot the accuracy relative to the accuracy when using the default size ($f_{\text{size}}=1.0$).
+Size Bias.
+Finally, we evaluate the impact of different sized foreground objects on the accuracy.
 For this evaluation, we use the mean foreground size strategy.
 We introduce a size factor $f_{\text{size}}$ by which we additionally scale the foreground object before pasting it onto the background.
-Results are again normalized by the accuracy when using the mean foreground size ($f_{\text{size}}=1.0$).
-Figure 4 shows the size bias curves of ViT-S and ViT-B when trained on ImageNet and ForNet.
-Models trained on ForNet maintain better performance even with smaller foreground objects, when ImageNet-trained models exhibit a more rapid performance decline.
-Therefore, ForNet-training improves robustness to variations in object scale.
+Results are normalized by the accuracy when using $f_{\text{size}}=1.0$.
+Figure 5 shows the size bias curves of models trained with and without ForAug.
+Models trained using ForAug maintain perform better, especially with smaller foreground objects.
+Therefore, ForAug-training improves robustness to variations in object scale, especially for larger models.
 5 Discussion & Conclusion
 We introduce ForAug, a novel data augmentation scheme that facilitates improved Transformer training for image classification.
-By explicitly separating and recombining foreground objects and backgrounds, ForAug enables controlled data augmentation, leading to significant performance gains on ImageNet and downstream fine-grained classification tasks.
+By explicitly separating and recombining foreground objects and backgrounds, ForAug enables controlled data augmentation beyond existing image compositions, leading to significant performance gains on ImageNet and downstream fine-grained classification tasks.
 Furthermore, ForAug provides a powerful framework for analyzing model behavior and quantifying biases, including background robustness, foreground focus, center bias, and size bias.
-Our experiments demonstrate that training on ForNet, the instantiation of ForAug on ImageNet, not only boosts accuracy but also significantly reduces these biases, resulting in more robust and generalizable models.
+Our experiments demonstrate that training using ForAug not only boosts accuracy but also significantly reduces these biases, resulting in more robust and generalizable models.
 In the future, we see ForAug be also applied to other datasets and tasks, like video recognition or segmentation.
@@
 This work was funded by the Carl-Zeiss Foundation under the Sustainable Embedded AI project (P2021-02-009) and by the EU project SustainML (Horizon Europe grant agreement No 101070408).
-All compute was done thanks to the Pegasus cluster at DFKI.
-Appendix A Training Setup
+All compute was done thanks to the Pegasus cluster at DFKI Kaiserslautern.
+Appendix A Extended Bates Distribution
+Figure 6: Plot of the probability distribution function (PDF) of the extended Bates distribution for different parameters $\eta$.
+Higher values of $\eta$ concentrate the distribution around the center.
+We introduce an extension of the Bates distribution [1] to include negative parameters, enabling sampling of foreground object positions away from the image center.
+The standard Bates distribution, for $\eta\in\mathbb{N}$, is defined as the mean of $\eta$ independent random variables drawn from a uniform distribution [19].
+A larger $\eta$ value increases the concentration of samples around the distribution’s mean, which in this case is the image center.
+To achieve an opposite effect–concentrating samples at the image borders–we extend the distribution to $\eta\leq 1$.
+$X\sim\text{Bates}(\eta):\Leftrightarrow s(X)\sim\text{Bates}(-\eta)$
+This is accomplished by sampling from a standard Bates distribution with parameter $\eta\geq 1$ and then applying a sawtooth function. The sawtooth function on the interval $[0,1]$ is defined as
+$s(x)=\begin{cases}x+0.5&\text{if }0<x<0.5\\ x-0.5&\text{if }0.5\leq x\leq 1\end{cases}$ (4)
+This function effectively maps the central portion of the interval to the edges and the edge portions to the center.
+For example, a value of 0.3 (central-left) is mapped to 0.8 (edge-right), while 0.8 (edge-right) is mapped to 0.3 (central-left).
+This transformation inverts the distribution’s concentration, shifting the probability mass from the center to the borders.
+We visualize the distribution function of the extended Bates distribution in Figure 6.
+Both $\eta=1$ and $\eta=-1$ result in a uniform distribution across the image.
+Appendix B Resource Usage of ForAug
+To utilize the proposed ForAug, specific computational resources are necessary, particularly for computing and storing for the output of the segmentation stage and for on-the-fly processing of the recombination stage.
+Segmentation.
+ForAug involves a computationally expensive segmentation and infill stage, which is a one-time calculation per dataset.
+Once computed, the segmentation and infill results can be perpetually reused, amortizing the initial cost over all subsequent experiments and applications.
+On NVIDIA H100 GPUs, the segmentation stage will compute at a rate of $374.3\frac{\text{img}}{\text{GPU}\times\text{h}}$ when using Attentive Eraser or $5338.6\frac{\text{img}}{\text{GPU}\times\text{h}}$ for LaMa.
+For ImageNet this comes down to just under 9 days (Attentive Eraser) or 16 hours (LaMa) on two 8 GPU nodes.
+To facilitate immediate use and reproduction of results, we publicly provide the precalculated segmentation stage output for the ImageNet dataset for download [^2] Link will go here..
+The output of ForAug’s segmentation step on ImageNet dataset requires 73 GB of additional disk space for the segmentation output, which is separate from the base 147 GB ImageNet size.
+Recombination.
+The recombination step of ForAug is implemented as a based data loader operation.
+It’s thus offloaded to the CPU, where it can be heavily parallelized and thus only results in a very minor increase in the training step-time.
+For example, using a ViT-B model on an NVIDIA A100 GPU, the average update step-time increased by $1\%$, from $528\pm 2$ ms to $534\pm 1$ ms.
+Appendix C Training Setup
+Table 9: Training setup and hyperparameters for our ImageNet training.
+Parameter ViT, Swin, ResNet DeiT
+Image Resolution $\times 224 224×224$$\times 224$
+Epochs 300 300
+Learning Rate 3e-3 S/B: 1e-3, L: 5e-4
+Learning Rate Schedule cosine decay cosine decay
+Batch Size 2048 1024
+GPUs $4\times$ NVIDIA A100/H100/H200 $4\times$ NVIDIA A100/H100/H200
+Warmup Schedule linear linear
+Warmup Epochs 3 3
+Weight Decay 0.02 0.05
+Label Smoothing 0.1 0.1
+Optimizer Lamb [59] AdamW
+Data Augmentation Policy 3-Augment [48] DeiT [47]
+Augmentations
+Resize
+RandomCrop
+HorizontalFlip
+Grayscale
+Solarize
+GaussianBlur
+ColorJitter
+CutMix [61]
+RandomResizedCrop
+HorizontalFlip
+RandomErase [64]
+RandAugment [5]
+ColorJitter
+Mixup [63]
+CutMix [61]
+Table 10: Training setup for finetuning on different downstream datasets.
+Other settings are the same as in Table 9.
+For finetuning, we always utilize 3-Augment and the related parameters from the ViT, Swin, ResNet column of Table 9
+Dataset Batch Size Epochs Learning Rate Num. GPUs
+Aircraft 512 500 3e-4 2
+Cars 1024 500 3e-4 4
+Flowers 256 500 3e-4 1
+Food 2048 100 3e-4 4
+Pets 512 500 3e-4 2
+On ImageNet we use the same training setup as [32] and [48] without pretraining for ViT, Swin, and ResNet.
+For DeiT, we train the same ViT architecture but using the data augmentation scheme and hyperparameters from [47].
+As our focus is on evaluating the changes in accuracy due to ForAug, like [32], we stick to one set of hyperparameters for all models.
+We list the settings used for training on ImageNet in Table 9 and the ones used for finetuning those weights on the downstream datasets in Table 10.
+Out implementation is using PyTorch [35] and the timm library [55] for model architectures and basic functions.
+Table 11: Hardware and Software specifics used for both training and evaluation.
 Parameter Value
-Image Resolution $224\times 224$
-Epochs 300
-Learning Rate 3e-3
-Learning Rate Schedule cosine decay
-Batch Size 2048
-Warmup Schedule linear
-Warmup Epochs 3
-Weight Decay 0.02
-Label Smoothing 0.1
-Optimizer Lamb [57]
-Data Augmentation Policy 3-Augment [48]
-Table 9: Training setup for our ImageNet and ForNet training.
-Dataset Batch Size Epochs Learning Rate
-Aircraft 512 500 3e-4
-Cars 1024 500 3e-4
-Flowers 256 500 3e-4
-Food 2048 100 3e-4
-Pets 512 500 3e-4
-Table 10: Training setup for finetuning on different downstream datasets. Other settings are the same as in Table 9.
-On ImageNet we use the same training setup as [33] and [48] without pretraining.
-As our focus is on evaluating the changes in accuracy due to ForAug/ForNet, like [33], we stick to one set of hyperparameters for all models.
-We list the settings used for training on ImageNet and ForNet in Table 9 and the ones used for finetuning those weights on the downstream datasets in Table 10.
-Appendix B Infill Model Comparison
+GPU NVIDIA A100/H100/H200
+CPU 24 CPU cores (Intex Xenon) per GPU
+Memory up to 120GB per GPU
+Operating System Enroot container for SLURM based on Ubuntu 24.04 LTS
+Python 3.12.3
+PyTorch 2.7.0
+TorchVision 0.22.0
+Timm 1.0.15
+Table 11 lists the specific hardware we use, as well as versions of the relevant software packages.
+Appendix D Infill Model Comparison
+Table 12: Example infills of LaMa and Attentive Eraser.
 LaMa Att. Eraser LaMa Att. Eraser
@@
 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 11: Example infills of LaMa and Attentive Eraser.
-Appendix C Images with High Infill Ratio
+We visualize example infilled images for both LaMa [45] and Attentive Eraser [43] in Table 12.
+We qualitatively find that while LaMa often leaves repeated textures of blurry spots where the object was erased, Attentive Eraser produces slightly cleaner and more coherent infills of the background.
+Appendix E Image Infill Ratio
+Table 13: Example infills with a large relative foreground area size that is infilled (infill ratio).
 Infill Ratio LaMa Att. Eraser
@@
 88.2 [Uncaptioned image] [Uncaptioned image]
-Table 12: Example infills with a large relative foreground area size that is infilled (infill ratio).
+Figure 7: We plot the distribution of the relative size of the detected foreground object that is infilled in our Segmentation step of ImageNet.
+While most images contain objects of smaller size, there is a peak where Grounded SAM [37] detects almost the whole image as the foreground object.
+For examples of such large infills, see Table 13.
+Table 13 shows infills for images where Grounded SAM [37] marks a high percentile of the image as the foreground object (Infill Ratio), that has to be erased by the infill models.
+While LaMa tends to fill those spots with mostly black or gray and textures similar to what we saw in Table 12, Attentive Eraser tends to create novel patterns by copying what is left of the background all over the rest of the image.
+Figure 7 plots the distribution of infill ratios in ForAug.
+While there is a smooth curve of the number of detections decreasing with the infill ratio until $\approx 90\%$, there is an additional peak at $\approx 100\%$ infill ratio.
+We believe that this peak is made up of failure cases of Grounded SAM.
+We filter out all backgrounds that have an infill ratio larger than our pruning threshold $t_{\text{prune}}=0.8$, which translates to $10\%$ of backgrounds.